### v2

# Infinite-time Mean Field FBSDEs and Viscosity Solutions to Elliptic Master Equations

Abstract
This paper presents a further investigation of the properties of infinite-time mean field FBSDEs and elliptic master equations, which were introduced in [14] as mathematical tools for solving discounted infinite-time mean field games. By establishing the continuous dependence of the FBSDE solutions on their initial values, we prove the flow property of the mean field FBSDEs. Furthermore, we prove that, at the Nash equilibrium, the value function of the representative player constitutes a viscosity solution to the corresponding elliptic master equation. Our work extends the classical theory of finite-time mean field games and parabolic master equations to the infinite-time setting.
Keywords. discounted infinite-time mean field games, infinite-time FBSDEs, elliptic master equations, viscosity solution.
1 Introduction
The study of mean field games was initiated independently by Lasry-Lions (see [7, 8, 9]) and Huang-Malhamé-Caines [6], which is an analysis of limit models for symmetric weakly interacting $(N+1)$-player differential games. The parabolic master equation plays a crucial role in the analysis of mean field games, which was introduced by Lions in lectures [10]. It describes the strategic interaction between a representative player and the collective environment. We refer the reader to [4, 3, 5] for a comprehensive exposition on the subject.
Forward-backward stochastic differential equations (FBSDEs) also serve as a powerful tool for the study of mean field games. The investigation of general nonlinear BSDEs was pioneered by Pardoux and Peng [11, 12] in the early 1990s. [13] studied the infinite-time FBSDEs and established connections with quasilinear elliptic PDEs. Recently, [1] extended this framework to the Mckean-Vlasov FBSDEs. This paper establishes the existence and uniqueness theorem for a broader class of infinite-time FBSDEs, and employs it to address the existence of viscosity solutions for elliptic master equations.
In the recent work [14], we proposed the discounted infinite-time mean field game model and elliptic master equation, which extends the traditional framework to infinite-time case. Within this framework, we introduced the following two systems of infinite-time forward-backward stochastic differential equations (FBSDEs). The first one represents the state process of the social equilibrium, while the second denotes the state process of the representative player with initial state $x$.
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\partial_{y}H(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[\partial_{x}{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})-rY_{t}^{\xi}\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi,\end{cases}$ (1.1)
$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.2)
Here $r>0$ is the discount factor and
$H(x,\mu,y)\triangleq\min_{a\in\mathbb{R}}\left[b(x,\mu,a)\cdot y+f(x,\mu,a)\right],$ (1.3)
and $\hat{\alpha}(x,y)$ is the unique minimizer. Because we assume that $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$ and $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$, the value of minimizer $\hat{\alpha}$ is independent of $\mu$. After further assuming that $f$ and $b$ possess good smoothness properties, we can obtain the following relationship:
$\partial_{y}H(x,\mu,y)=b(x,\mu,\hat{\alpha}(x,y)).$ (1.4)
In [14], we introduced the elliptic master equation:
$\begin{split}rU(x,\mu)=&H(x,\mu,\partial_{x}U(x,\mu))+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (1.5)
Here $\partial_{x},\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu},\partial_{\tilde{x}\mu}$ are $W_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law. We have proved that if the master equation (1.5) admits a classical solution with sufficient regularity, we derive the following representation for Equation (1.1) and (1.2):
$Y_{t}^{\xi}=\partial_{x}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{\xi}=\partial_{xx}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),$ (1.6)
$Y_{t}^{x}=\partial_{x}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x}=\partial_{xx}U(X_{t}^{x},\mathcal{L}_{X_{t}^{\xi}}).$ (1.7)
In this paper, we prove that (1.1) and (1.2) possess the flow property:
$X_{t}^{x,\xi}|_{x=\xi}=X_{t}^{\xi},\quad Y_{t}^{x,\xi}|_{x=\xi}=Y_{t}^{\xi},\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$},$ (1.8)
and
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (1.9)
Furthermore, we prove that the value function of the representative player
$V(x,\mu)=\mathbb{E}\bigg[\int_{0}^{+\infty}\mathrm{e}^{-rt}f\big(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x,\xi},Y_{t}^{x,\xi})\big)\mathrm{d}t\bigg]$ (1.10)
is a viscosity solution to the master equation (1.5).
2 Preliminaries
We will use the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$. Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$, such that it can support any probability measure on $\mathbb{R}$ with finite second moment.
Let $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ be a copy of the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with corresponding Brownian motion $B^{\prime}$, define the larger filtered probability space by
$\tilde{\Omega}\triangleq\Omega\times\Omega^{\prime},\quad\tilde{\mathcal{F}}\triangleq\mathcal{F}\otimes\mathcal{F}^{\prime}\quad\tilde{\mathbb{F}}=\{\tilde{\cal F}_{t}\}_{t\geq 0}\triangleq\{{\cal F}_{t}\otimes{\cal F}^{\prime}_{t}\}_{t\geq 0},\quad\tilde{\mathbb{P}}\triangleq\mathbb{P}\otimes\mathbb{P}^{\prime},\quad\tilde{\mathbb{E}}\triangleq\mathbb{E}^{\tilde{\mathbb{P}}}$. (2.1)
[…]
Let ${\cal P}\triangleq{\cal P}(\mathbb{R})$ be the set of all probability measures on $\mathbb{R}$ and let ${\cal P}_{p}(p\geq 1)$ denote the set of $\mu\in{\cal P}$ with finite $p$-th moment. For any sub-$\sigma$-field $\cal G$$\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, $\cal G$}-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ . For any $\mu,\nu\in{\cal P}_{p}$, we define the ${\cal W}_{p}$–Wasserstein distance between them as follows:
${\cal W}_{p}(\mu,\nu):=\inf\Big\{\big(\mathbb{E}[|\xi-\eta|^{q}]\big)^{1/q}:~\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big\}.$
Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}$-valued adapted stochastic process ($v_{t}$) start from $t_{0}$ such that
$\mathbb{E}\left[\int_{t_{0}}^{\infty}e^{-Kt}|v_{t}|^{2}dt\right]<+\infty.$ (2.2)
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$.
We introduce the Wasserstein space and differential calculus on Wasserstein space. For a ${\cal W}_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$, its ${\cal W}_{2}$-Wasserstein derivatives [4](also called Lions-derivative), takes the form $\partial_{\mu}U:(\mu,\tilde{x})\in{\cal P}_{2}\times\mathbb{R}\to\mathbb{R}$ and satisfies:
$U({\cal L}_{\xi+\eta})-U(\mu)=\mathbb{E}\big[\langle\partial_{\mu}U(\mu,\xi),\eta\rangle\big]+o(\|\eta\|_{2}),\ \forall\ \xi\in\mathbb{L}^{2}(\mathcal{F};\mu),\eta\in\mathbb{L}^{2}(\mathcal{F}).$ (2.3)
Let ${\cal C}^{0}({\cal P}_{2})$ denote the set of ${\cal W}_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$. For ${\cal C}^{1}({\cal P}_{2})$, we mean the space of functions $U\in{\cal C}^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.3). Let ${\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$. Let ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U,\partial_{xxx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U,\partial_{xx\mu}U,\partial_{x\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula [2, 4]. For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$. Fix $T>0$ and let all conditions be restricted to the interval [0,T]. Then we have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big[\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\partial_{xx}U[\sigma_{t}^{1}]^{2}\Big](t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big(\tilde{\mathbb{E}}_{{\cal F}_{t}}\big[\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})[\tilde{\sigma}_{t}^{2}]^{2}\big]\Big)\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.4)
Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ .
3 Solutions to infinite-time FBSDEs
For the needs of subsequent problems, we aim to establish a more general theorem on the existence and uniqueness of solutions for infinite-time FBSDEs. Consider the following form of infinite-time FBSDEs:
$\begin{cases}\mathrm{d}X_{t}=G(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,\omega,X_{t},Y_{t},\mathcal{L}_{X_{t}})\mathrm{d}t+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi,\end{cases}$ (3.1)
where $G,F:\mathbb{R}_{+}\times\Omega\times\mathbb{R}\times\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ are two progressively measurable functions, $\sigma\in\mathbb{R}$ is a constant and $\xi$ is an ${\cal F}_{0}$-measurable square integrable random variable. For any $(v_{t})\in L_{K}^{2}$, we define the exponentially weighted $L^{2}$ norm
$\|v\|_{K}^{2}\triangleq\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-Kt}|v_{t}|^{2}\mathrm{d}t\right].$ (3.2)
For simplicity, we only solve (3.1) for one dimensional $(X_{t},Y_{t},Z_{t})$ and starting time $t_{0}=0$ , but our result can be easily generalized to multidimensional case and arbitrary starting time $t_{0}>0$. The key idea of our proof follows [1, 13].
Assumption 3.1
Assume that for some constant $K$, the functions $F$ and $G$ satisfy:
(i) For any $L_{K}^{2}$ processes $(X_{t},Y_{t})$, $G(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})$ and $F(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})$ belong to $L_{K}^{2}$.
(ii) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
$|G(t,\omega,x,y,\mu)-G(t,\omega,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,\omega,x,y,\mu)-F(t,\omega,x^{\prime},y^{\prime},\mu^{\prime})|$ (3.3)
$\leq$ $\ell(|x-x^{\prime}|+|y-y^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime})).\quad\mathrm{a.s.}$
(iii) There exists a constant $\kappa>K/2$, such that for any $t\geq 0$ and any square integrable random variables $X,X^{\prime},Y,Y^{\prime}$,
$\mathbb{E}\left[-K\hat{X}\hat{Y}-\hat{X}(F(t,\omega,U)-F(t,\omega,U^{\prime}))+\hat{Y}(G(t,\omega,U)-G(t,\omega,U^{\prime}))\right]$ (3.4)
$\leq-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}\triangleq X-X^{\prime},\hat{Y}\triangleq Y-Y^{\prime}$ and $U\triangleq(X,Y,\mathcal{L}_{X}),U^{\prime}\triangleq(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}}).$
Theorem 3.2
Under Assumption 3.1, for each ${\cal F}_{0}$-measurable square integrable random variable $\xi$ , (3.1) has a unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{K}^{2}$.
Proof.  First, we prove the uniqueness. Suppose there exist two solutions $(X_{t},Y_{t},Z_{t})$, $(X_{t}^{\prime},Y_{t}^{\prime},Z_{t}^{\prime})$ in $L_{K}^{2}$ to (3.1), and denote
$\hat{X}\triangleq X-X^{\prime}\quad\hat{Y}\triangleq Y-Y^{\prime}\quad\hat{Z}\triangleq Z-Z^{\prime}.$ (3.5)
We choose a sequence of $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-KT_{i}}\hat{X}_{T_{i}}\hat{Y}_{T_{i}}\right]\rightarrow 0.$ (3.6)
Applying Itô’s formula to $\mathrm{e}^{-Kt}\hat{X}_{t}\hat{Y}_{t}$, we get that
$\mathbb{E}\left[\mathrm{e}^{-KT_{i}}\hat{X}_{T_{i}}\hat{Y}_{T_{i}}\right]$ (3.7)
[…]
$\left.\left.+\hat{Y}_{t}(G(t,\omega,X_{t},Y_{t},{\cal L}_{X_{t}})-G(t,\omega,X_{t}^{\prime},Y_{t}^{\prime},{\cal L}_{X_{t}^{\prime}}))\right)dt\right]$
$\leq$ $-\kappa\mathbb{E}\left[\int_{0}^{T_{i}}e^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right].$
Letting $T_{i}\to\infty$, we get that
$\|\hat{X}\|_{K}^{2}=\|\hat{Y}\|_{K}^{2}=0,$ (3.8)
and hence we complete the proof of the uniqueness.
Next, we prove the existence of solutions, for this purpose, we use the continuity method. We study the following family of infinite-time FBSDEs parametrized by $\lambda\in[0,1]$,
$\begin{cases}\begin{aligned} \mathrm{d}X_{t}^{\lambda}=&\left[\lambda G(t,\omega,X_{t}^{\lambda},Y_{t}^{\lambda},\mathcal{L}_{X_{t}^{\lambda}})\right.\left.-\kappa(1-\lambda)Y_{t}^{\lambda}+\phi_{t}(\omega)\right]\mathrm{d}t+\sigma\mathrm{d}B_{t},\end{aligned}\\ \begin{aligned} \mathrm{d}Y_{t}^{\lambda}=&-\left[\lambda F(t,\omega,X_{t}^{\lambda},Y_{t}^{\lambda},\mathcal{L}_{X_{t}^{\lambda}})\right.\left.+\kappa(1-\lambda)X_{t}^{\lambda}+\psi_{t}(\omega)\right]\mathrm{d}t+Z_{t}^{\lambda}\mathrm{d}B_{t},\end{aligned}\\ \begin{aligned} &X_{0}^{\lambda}=\xi,\quad(X_{t}^{\lambda},Y_{t}^{\lambda},Z_{t}^{\lambda})\in L_{K}^{2},\end{aligned}\end{cases}$ (3.9)
where $\phi,\psi$ are two arbitrary processes in $L_{K}^{2}$. Note that when $\lambda=1,\phi=\psi=0$, (3.9) becomes (3.1), and when $\lambda=0$, (3.9) becomes
$\begin{cases}&\mathrm{d}X_{t}^{0}=(-\kappa Y_{t}^{0}+\phi_{t}(\omega))\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ &\mathrm{d}Y_{t}^{0}=-(\kappa X_{t}^{0}+\psi_{t}(\omega))\mathrm{d}t+Z_{t}^{0}\mathrm{d}B_{t},\\ &X_{0}^{0}=\xi.\end{cases}$ (3.10)
It has been proved in ([1], Lemma 2.1) that (3.10) has unique solution $(X^{0},Y^{0},Z^{0})\in L_{K}^{2}$.
Now, suppose for some $\lambda_{0}\in[0,1)$, we have that for any ${\cal F}_{0}$-measurable square integrable random variable $\xi$ and $\phi,\psi\in L^{2}_{K}$, (3.9) has a unique solution $(X^{\lambda_{0}},Y^{\lambda_{0}},Z^{\lambda_{0}})$ in $L^{2}_{K}$. We try to find a constant $\delta_{0}$, such that for any $\delta\in[0,\delta_{0}]$, the FBSDE(3.9) has a unique solution $(X^{\lambda_{0}+\delta},Y^{\lambda_{0}+\delta},Z^{\lambda_{0}+\delta})$ in $L^{2}_{K}$ for any given $\xi,\phi,\psi$. To do this, we consider the following FBSDE:
$\begin{cases}\begin{aligned} \mathrm{d}X_{t}={}&\bigl[\lambda_{0}G(t,\omega,X_{t},Y_{t},\mathcal{L}_{X_{t}})-\kappa(1-\lambda_{0})Y_{t}\\ &+\delta\bigl(G(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})+\kappa y_{t}\bigr)+\phi_{t}(\omega)\bigr]\mathrm{d}t+\sigma\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}={}&-\bigl[\lambda_{0}F(t,\omega,X_{t},Y_{t},\mathcal{L}_{X_{t}})+\kappa(1-\lambda_{0})X_{t}\\ &+\delta\bigl(F(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})-\kappa x_{t}\bigr)+\psi_{t}(\omega)\bigr]\mathrm{d}t+Z_{t}\mathrm{d}B_{t},\\ X_{0}={}&\xi.\end{aligned}\end{cases}$ (3.11)
For any pair $(x_{t},y_{t})$ in $L_{K}^{2}$, we have $G(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})$ and $F(t,\omega,x_{t},y_{t},\mathcal{L}_{x_{t}})$ belong to $L_{K}^{2}$ according to our hypothesis, then the FBSDE (3.11) admits a unique solution $(X,Y,Z)$ in $L_{K}^{2}$. We can define a map $\Phi$ through
$\Phi:(x,y)\mapsto(X,Y).$ (3.12)
We then prove that the map $\Phi$ is a contraction on $L_{K}^{2}$.
Take another pair $(x_{t}^{\prime},y_{t}^{\prime})$ in $L_{K}^{2}$ and its image $(X_{t}^{\prime},Y_{t}^{\prime})$. Denote
$U_{t}=(X_{t},Y_{t},{\cal L}_{X_{t}}),\quad u_{t}=(x_{t},y_{t},{\cal L}_{x_{t}}),$ (3.13)
$\hat{X}_{t}=X_{t}-X_{t}^{\prime},\quad\hat{Y}_{t}=Y_{t}-Y_{t}^{\prime},$
$\hat{x}_{t}=x_{t}-x_{t}^{\prime},\quad\hat{y}_{t}=y_{t}-y_{t}^{\prime}.$
Applying Itô’s formula to $\mathrm{e}^{-Kt}\hat{X}_{t}\hat{Y}_{t}$, we get that
$\left[\mathrm{e}^{-KT}\hat{X}_{T}\hat{Y}_{T}\right]$ (3.14)
$=\lambda_{0}\mathbb{E}\biggl[\int_{0}^{T}\mathrm{e}^{-Kt}\biggl(-K\hat{X}_{t}\hat{Y}_{t}-\hat{X}_{t}\bigl(F(t,\omega,U_{t})-F(t,\omega,U_{t}^{\prime})\bigr)\quad+\hat{Y}_{t}\bigl(G(t,\omega,U_{t})-G(t,\omega,U_{t}^{\prime})\bigr)\biggr)\mathrm{d}t\biggr]$
$-\kappa(1-\lambda_{0})\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right]$
$-(K-\lambda_{0}K)\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\hat{X}_{t}\hat{Y}_{t}\mathrm{d}t\right]$
$+\kappa\delta\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{X}_{t}\hat{x}_{t}+\hat{Y}_{t}\hat{y}_{t}\right)\mathrm{d}t\right]$
$+\delta\mathbb{E}\biggl[\int_{0}^{T}\mathrm{e}^{-Kt}\biggl(-\hat{X}_{t}\bigl(F(t,\omega,u_{t})-F(t,\omega,u_{t}^{\prime})\bigr)\quad+\hat{Y}_{t}\bigl(G(t,\omega,u_{t})-G(t,\omega,u_{t}^{\prime})\bigr)\biggr)\mathrm{d}t\biggr].$
It holds that
$\mathbb{E}\biggl[-K\hat{X}_{t}\hat{Y}_{t}-\hat{X}_{t}\bigl(F(t,\omega,U_{t})-F(t,\omega,U_{t}^{\prime})\bigr)+\hat{Y}_{t}\bigl(G(t,\omega,U_{t})-G(t,\omega,U_{t}^{\prime})\bigr)\biggr]$ (3.15)
$\leq-\kappa\mathbb{E}\left[\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right]$
and
$\mathbb{E}\left[-\hat{X}_{t}\bigl(F(t,\omega,u_{t})-F(t,\omega,u_{t}^{\prime})\bigr)+\hat{Y}_{t}\bigl(G(t,\omega,u_{t})-G(t,\omega,u_{t}^{\prime})\bigr)\right]$ (3.16)
$\leq$ $\mathbb{E}\left[|\hat{X}_{t}|\cdot|F(t,\omega,u_{t})-F(t,\omega,u_{t}^{\prime})|+|\hat{Y}_{t}|\cdot|G(t,\omega,u_{t})-G(t,\omega,u_{t}^{\prime})|\right]$
$\leq$ $\frac{3\ell}{2}\mathbb{E}\left[|\hat{X}_{t}|^{2}+|\hat{Y}_{t}|^{2}\right]+2\ell\mathbb{E}\left[|\hat{x}_{t}|^{2}+|\hat{y}_{t}|^{2}\right].$
So we have
$\mathbb{E}\left[\mathrm{e}^{-KT}\hat{X}_{T}\hat{Y}_{T}\right]\leq$ $-\left(\kappa-\frac{K}{2}-\frac{\kappa\delta+3l\delta}{2}\right)\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right]$ (3.17)
$+\frac{\kappa\delta+4l\delta}{2}\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-Kt}\left(\hat{x}_{t}^{2}+\hat{y}_{t}^{2}\right)\mathrm{d}t\right].$
We take
$\delta_{0}=\frac{2\kappa-K}{3\kappa+11\ell}$ (3.18)
and choose a sequence of $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-KT_{i}}\hat{X}_{T_{i}}\hat{Y}_{T_{i}}\right]\rightarrow 0.$ (3.19)
For any $\delta\in[0,\delta_{0}]$, we have
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-Kt}\left(\hat{X}_{t}^{2}+\hat{Y}_{t}^{2}\right)\mathrm{d}t\right]\leq\frac{1}{2}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-Kt}\left(\hat{x}_{t}^{2}+\hat{y}_{t}^{2}\right)\mathrm{d}t\right].$ (3.20)
Therefore $\Phi$ is a contraction.
By repeating this procedure for $1/\delta_{0}]$ many times, we conclude that there exists a solution to (3.9) with $\lambda=1$. In particular, we get a $L_{K}^{2}$ solution to (3.1).
4 Properties of mean field FBSDEs
In this section, we investigate the following mean-field FBSDEs:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\partial_{y}H(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[\partial_{x}{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})-rY_{t}^{\xi}\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi,\end{cases}$ (4.1)
$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (4.2)
To obtain further properties of the above two equations, we require the following assumptions:
Assumption 4.1
(i) $H(x,\mu,y)$ is jointly continuous and and all second-order partial derivatives exist.
(ii) $\partial_{y}H(x,\mu,y)$ and $\partial_{x}H(x,\mu,y)$ are Lipschitz continuous in $(x,\mu,y)$. More specifically, there exists a constant $\ell>0$ such that
$|\partial_{y}H(x,\mu,y)-\partial_{y}H(x^{\prime},\mu^{\prime},y^{\prime})|\leq\ell\left(|x-x^{\prime}|+|y-y^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})\right),$ (4.3)
$|\partial_{x}H(x,\mu,y)-\partial_{x}H(x^{\prime},\mu^{\prime},y^{\prime})|\leq\ell\left(|x-x^{\prime}|+|y-y^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})\right).$
(iii)There exist constants $\kappa,C_{0}>0$ , such that $C_{0}+r/2<\kappa$ and
$-(x-x^{\prime})\left[\partial_{x}H(x,\mu,y)-\partial_{x}H(x^{\prime},\mu^{\prime},y^{\prime})\right]+(y-y^{\prime})\left[\partial_{y}H(x,\mu,y)-\partial_{y}H(x^{\prime},\mu^{\prime},y^{\prime})\right]$ (4.4)
$\leq-\kappa\left(|x-x^{\prime}|^{2}+|y-y^{\prime}|^{2}\right)+C_{0}{\cal W}_{2}^{2}(\mu,\mu^{\prime}).$
Theorem 4.2
Under Assumption 4.1, both FBSDE (4.1) and (4.2) have unique solutions in $L_{r}^{2}$.
Proof.  For FBSDE (4.1), it’s clear that the conditions in Assumption 3.1 (i) and (ii) are satisfied. Taking four arbitrary square integrable random variables $X,X^{\prime},Y,Y^{\prime}$, we have
$\mathbb{E}\biggl[$ $-r\hat{X}\cdot\hat{Y}$ (4.5)
$-\hat{X}\left[\partial_{x}H(X,{\cal L}_{X},Y)-\partial_{x}H(X^{\prime},{\cal L}_{X^{\prime}},Y^{\prime})-r\hat{Y}\right$]
$+\hat{Y}\left[\partial_{y}H(X,{\cal L}_{X},Y)-\partial_{y}H(X^{\prime},{\cal L}_{X^{\prime}},Y^{\prime})\right]\biggr]$
$\leq$ $-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right]+C_{0}\mathcal{W}_{2}^{2}({\cal L}_{X},{\cal L}_{X^{\prime}})$
$\leq$ $-(\kappa-C_{0})\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}\triangleq X-X^{\prime},\hat{Y}\triangleq Y-Y^{\prime}$. Since $\kappa-C_{0}>r/2$, the conditions in Assumption 3.1 (iii) is satisfied. Then FBSDE (4.1) has a unique solution in $L_{r}^{2}$.
After solving FBSDE (4.1), we substitute its solution ${\cal L}_{X_{t}^{\xi}}$ into FBSDE (4.2), and it can similarly be shown that there exists a unique solution to FBSDE (4.1).
The following proposition informs us that the solution of the mean field FBSDE exhibits favorable continuous dependence on the initial value, which is of paramount importance for our subsequent research.
Proposition 4.3
For FBSDEs (4.1) and (4.2), assuming all conditions in Assumption 4.1 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\xi,\xi^{\prime}\in\mathbb{L}^{2}({\cal F}_{0})$, there exists a constant $C>0$, such that
$\mathbb{E}\left[|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}|^{2}\right]+\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\leq C\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]$ (4.6)
and
$|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}|^{2}+\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\leq C\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (4.7)
Proof.  Set
$\begin{split}&\hat{X}^{\xi}=X^{\xi}-X^{\xi^{\prime}},\quad\hat{Y}^{\xi}=Y^{\xi}-Y^{\xi^{\prime}},\quad\hat{Z}^{\xi}=Z^{\xi}-Z^{\xi^{\prime}},\\ &\hat{X}^{x,\xi}=X^{x,\xi}-X^{x^{\prime},\xi^{\prime}},\quad\hat{Y}^{x,\xi}=Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}},\quad\hat{Z}^{x,\xi}=Z^{x,\xi}-Z^{x^{\prime},\xi^{\prime}}.\end{split}$ (4.8)
$C_{1},C_{2},C_{3},C_{4},C_{5},C_{6},C_{7}$ appeared in the following proof are all positive constants.
Applying Itô’s formula to $\mathrm{e}^{-rt}|Y^{\xi}_{t}-Y^{\xi^{\prime}}_{t}|^{2}$, and taking a sequence of $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-rt}|Y^{\xi}_{T_{i}}-Y^{\xi^{\prime}}_{T_{i}}|^{2}\right]\rightarrow 0,$ (4.9)
we get that
$\mathbb{E}\left[\left|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}\right|^{2}\right]=$ $\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[r\left|Y_{t}^{\xi}-Y_{t}^{\xi^{\prime}}\right|^{2}$ (4.10)
$\quad+2\hat{Y}^{\xi}\cdot\left(\partial_{x}H\left(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi}\right)-\partial_{x}H\left(X_{t}^{\xi^{\prime}},\mathcal{L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{\xi^{\prime}}\right)-r\hat{Y}^{\xi}\right)$
$\quad-\left|Z_{t}^{\xi}-Z_{t}^{\xi^{\prime}}\right|^{2}\biggr]\mathrm{d}t$
$\leq$ $C_{1}\left(\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\right).$
Applying Itô’s formula to $\mathrm{e}^{-rt}\hat{X}^{\xi}\hat{Y}^{\xi}$, we get
$-\mathbb{E}\left[\hat{X}^{\xi}_{0}\hat{Y}^{\xi}_{0}\right]=$ $\ \mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[-r\hat{X}^{\xi}_{t}\hat{Y}^{\xi}_{t}$ (4.11)
$+\hat{Y}^{\xi}_{t}\left(\partial_{y}H(X^{\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{\xi}_{t})-\partial_{y}H(X^{\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{\xi^{\prime}}_{t})\right)$
$-\hat{X}^{\xi}_{t}\left(\partial_{x}H(X^{\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{\xi}_{t})-\partial_{x}H(X^{\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{\xi^{\prime}}_{t})-r\hat{Y}^{\xi}_{t}\right)\biggr]\mathrm{d}t$
$\leq$ $-\frac{r}{2}\left(\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\right).$
Then we have
$\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\leq$ $\frac{2}{r}\mathbb{E}\left[\hat{X}^{\xi}_{0}\hat{Y}^{\xi}_{0}\right]$ (4.12)
$\leq$ $\frac{1}{2C_{1}}\mathbb{E}\left[\left|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}\right|^{2}\right]+\frac{32C_{1}}{r^{2}}\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right].$
Combining (4.10) and (4.12), we get that
$\mathbb{E}\left[\left|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}\right|^{2}\right]\leq C_{2}\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right].$ (4.13)
Substituting this into (4.12), we obtain
$\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\leq C_{3}\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right].$ (4.14)
Next, we employ the same approach to FBSDE (4.2). Applying Itô’s formula to $\mathrm{e}^{-rt}|Y^{x,\xi}_{t}-Y^{x^{\prime},\xi^{\prime}}_{t}|^{2}$, we get that
$\mathbb{E}\left[\left|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}\right|^{2}\right]=$ $\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[r\left|Y_{t}^{x,\xi}-Y_{t}^{x^{\prime},\xi^{\prime}}\right|^{2}$ (4.15)
$\quad+2\hat{Y}^{x,\xi}\cdot\left(\partial_{x}H\left(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}\right)-\partial_{x}H\left(X_{t}^{x^{\prime},\xi^{\prime}},\mathcal{L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}}\right)-r\hat{Y}^{x,\xi}\right)$
$\quad-\left|Z_{t}^{x,\xi}-Z_{t}^{x^{\prime},\xi^{\prime}}\right|^{2}\biggr]\mathrm{d}t$
$\leq$ $C_{4}\left(\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}\right).$
Applying Itô’s formula to $\mathrm{e}^{-rt}\hat{X}^{x,\xi}\hat{Y}^{x,\xi}$, we can get that
$-\mathbb{E}\left[\hat{X}^{x,\xi}_{0}\hat{Y}^{x,\xi}_{0}\right]=$ $\ \mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[-r\hat{X}^{x,\xi}_{t}\hat{Y}^{x,\xi}_{t}$ (4.16)
$+\hat{Y}^{x,\xi}_{t}\left(\partial_{y}H(X^{x,\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{x,\xi}_{t})-\partial_{y}H(X^{x^{\prime},\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{x^{\prime},\xi^{\prime}}_{t})\right)$
$-\hat{X}^{x,\xi}_{t}\left(\partial_{x}H(X^{x,\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{x,\xi}_{t})-\partial_{x}H(X^{x^{\prime},\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{x^{\prime},\xi^{\prime}}_{t})-r\hat{Y}^{x,\xi}_{t}\right)\biggr]\mathrm{d}t$
$\leq$ $-\frac{r}{2}\left(\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\right)+C_{0}\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}.$
Then we have
$\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\leq$ $\frac{1}{2C_{4}}\mathbb{E}\left[\left|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}\right|^{2}\right]+C_{5}\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (4.17)
Substituting it into (4.15), we have
$\mathbb{E}\left[\left|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}\right|^{2}\right]\leq$ $C_{6}\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (4.18)
At last we come back to (4.17), we get that
$\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\leq$ $C_{7}\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (4.19)
Now we complete the proof.
Benefiting from the continuous dependence on initial values of the solutions to the equations that we previously proved, we can demonstrate the flow property of the mean field FBSDEs. This precisely reflects the essence of the mean field game: the aggregation of innumerable representative players constitutes the equilibrium state of the system.
Theorem 4.4
Let Assumption 4.1 hold. For any $\xi\in\mathbb{L}^{2}({\cal F}_{0})$, we have
$X_{t}^{x,\xi}|_{x=\xi}=X_{t}^{\xi},\quad Y_{t}^{x,\xi}|_{x=\xi}=Y_{t}^{\xi},\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$},$ (4.20)
and
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (4.21)
Proof.  We prove this theorem in two steps.
Step 1. We first assume $\xi$ is discrete, that is
$\xi=\sum_{i=1}^{n}x_{i}I_{A_{i}},$ (4.22)
where $x_{i}\in\mathbb{R}$ are constants and $\{A_{i}\}\in{\cal F}_{0}$ is a partition of $\Omega$. For each $i\in\{1,2,\cdots,n\}$, we can solve the following FBSDE in $L_{r}^{2}$:
$\begin{cases}\mathrm{d}X_{t}^{x_{i},\xi}=\partial_{y}H(X_{t}^{x_{i},\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x_{i},\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x_{i},\xi}=-\left[\partial_{x}{H}(X_{t}^{x_{i},\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x_{i},\xi})-rY_{t}^{x_{i},\xi}\right]\mathrm{d}t+Z_{t}^{x_{i}}\mathrm{d}B_{t},\\ X_{0}^{x_{i}}=x_{i}.\end{cases}$ (4.23)
Now we define
$X_{t}(\omega)\triangleq\sum_{i=1}^{n}X_{t}^{x_{i},\xi}(\omega)\cdot I_{A_{i}}(\omega),$ (4.24)
and similarly define $Y$ and $Z$ in the same manner. Then $(X,Y,Z)\in L_{r}^{2}$.
Multiplying both sides of the individual integral equations by $I_{A_{i}}$ and summing over $i$, we derive the aggregated equations. For the forward process:
$\begin{split}X_{t}&=\sum_{i=1}^{n}X_{t}^{x_{i},\xi}I_{A_{i}}\\ &=\sum_{i=1}^{n}\left(x_{i}+\int_{0}^{t}\partial_{y}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})\mathrm{d}s+B_{t}\right)I_{A_{i}}\\ &=\sum_{i=1}^{n}x_{i}I_{A_{i}}+\int_{0}^{t}\sum_{i=1}^{n}\partial_{y}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})I_{A_{i}}\mathrm{d}s+\sum_{i=1}^{n}I_{A_{i}}B_{t}\\ &=\xi+\int_{0}^{t}\partial_{y}H(X_{s},\mathcal{L}_{X_{s}^{\xi}},Y_{s})\mathrm{d}s+B_{t}.\end{split}$ (4.25)
For the backward process:
$\begin{split}Y_{t}&=\sum_{i=1}^{n}Y_{t}^{x_{i},\xi}I_{A_{i}}\\ &=\sum_{i=1}^{n}\left(Y_{T}^{x_{i},\xi}+\int_{t}^{T}\left[\partial_{x}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})-rY_{s}^{x_{i},\xi}\right]\mathrm{d}s-\int_{t}^{T}Z_{s}^{x_{i},\xi}\mathrm{d}B_{s}\right)I_{A_{i}}\\ &=\sum_{i=1}^{n}Y_{T}^{x_{i},\xi}I_{A_{i}}+\int_{t}^{T}\sum_{i=1}^{n}\left[\partial_{x}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})-rY_{s}^{x_{i},\xi}\right]I_{A_{i}}\mathrm{d}s-\int_{t}^{T}\sum_{i=1}^{n}Z_{s}^{x_{i},\xi}I_{A_{i}}\mathrm{d}B_{s}\\ &=Y_{T}+\int_{t}^{T}\left[\partial_{x}H(X_{s},\mathcal{L}_{X_{s}^{\xi}},Y_{s})-rY_{s}\right]\mathrm{d}s-\int_{t}^{T}Z_{s}\mathrm{d}B_{s}.\end{split}$ (4.26)
Thus, the aggregated processes $(X,Y,Z)$ satisfy the following FBSDE:
$\begin{cases}\mathrm{d}X_{t}=\partial_{y}H(X_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-\left[\partial_{x}{H}(X_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t})-rY_{t}\right]\mathrm{d}t+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (4.27)
We have proved that, given $\mathcal{L}_{X_{t}^{\xi}}$, the above FBSDE has a unique solution in $L_{r}^{2}$. So
$\left\lVert X-X^{\xi}\right\rVert_{r}=\left\lVert Y-Y^{\xi}\right\rVert_{r}=0$ (4.28)
and
$\mathbb{E}\left[|Y_{0}-Y_{0}^{\xi}|^{2}\right]=0.$ (4.29)
Then we have
$X_{t}^{x,\xi}(\omega)|_{x=\xi(\omega)}=X_{t}^{\xi}(\omega),\quad Y_{t}^{x,\xi}(\omega)|_{x=\xi(\omega)}=Y_{t}^{\xi}(\omega),\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$},$ (4.30)
and
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (4.31)
Step 2. In the general case, let $\xi_{n}\in\mathbb{L}({\cal F}_{0})$ be a discrete approximation of $\xi$ such that $\mathbb{E}\left[\left\lvert\xi_{n}-\xi\right\rvert^{2}\right]\rightarrow 0$. By Step 1, we have
$X_{t}^{x,\xi_{n}}(\omega)|_{x=\xi_{n}(\omega)}=X_{t}^{\xi_{n}}(\omega),\quad Y_{t}^{x,\xi_{n}}(\omega)|_{x=\xi_{n}(\omega)}=Y_{t}^{\xi_{n}}(\omega)\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$}.$ (4.32)
Applying Proposition 4.3, we have
$\left\lVert X^{\xi_{n}}-X^{\xi}\right\rVert_{r}\to 0,\quad\left\lVert Y^{\xi_{n}}-Y^{\xi}\right\rVert_{r}\to 0.$ (4.33)
On the other hand,
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(X_{t}^{\xi(\omega),\xi}-X_{t}^{\xi_{n}(\omega),\xi_{n}}\right)^{2}\mathrm{d}t\right]\leq$ $\mathbb{E}\left[\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(X_{t}^{\xi(\omega),\xi}-X_{t}^{\xi_{n}(\omega),\xi_{n}}\right)^{2}\mathrm{d}t|{\cal F}_{0}\right]\right]$ (4.34)
$\leq$ $C\mathbb{E}\left[|\xi(\omega)-\xi_{n}(\omega)|^{2}+\mathbb{E}\left[|\xi-\xi_{n}|^{2}\right]\right]\rightarrow 0.$
So we have $X_{t}^{\xi}$ and $X_{t}^{x,\xi}|_{x=\xi}$ are the same in $L_{r}^{2}$, The same holds for $Y_{t}^{\xi}$ and $Y_{t}^{x,\xi}|_{x=\xi}$. Then we conclude that
$X_{t}^{x,\xi}|_{x=\xi}=X_{t}^{\xi},\quad Y_{t}^{x,\xi}|_{x=\xi}=Y_{t}^{\xi},\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$}.$ (4.35)
In addition, by Proposition 4.3, we also have
$Y_{0}^{x,\xi_{n}}|_{x=\xi_{n}}=Y_{0}^{\xi_{n}},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (4.36)
and
$\mathbb{E}\left[|Y_{0}^{\xi}-Y_{0}^{\xi_{n}}|^{2}\right]\leq C\mathbb{E}\left[|\xi-\xi_{n}|^{2}\right]\to 0,$ (4.37)
$\mathbb{E}\left[|Y_{0}^{\xi(\omega),\xi}-Y_{0}^{\xi_{n}(\omega),\xi_{n}}|^{2}\right]=\mathbb{E}\left[\mathbb{E}\left[|Y_{0}^{\xi(\omega),\xi}-Y_{0}^{\xi_{n}(\omega),\xi_{n}}|^{2}|{{\cal F}_{0}}\right]\right]$
$\leq C\mathbb{E}\left[|\xi(\omega)-\xi_{n}(\omega)|^{2}+\mathbb{E}\left[|\xi-\xi_{n}|^{2}\right]\right]\rightarrow 0.$
So we have
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (4.38)
5 Viscosity solution to master equation
In this section, we will prove the main result of this paper: the value function (1.10) is the viscosity solution to the master equation (1.5).
5.1 Differentiability of mean-field FBSDEs
In order to prove the main theorem, we first establish the differentiability of the solutions to the mean field FBSDEs. For this purpose, we require the following assumptions:
Assumption 5.1
(i) $H(x,\mu,y)$ has at most quadratic growth.
(ii) There exist constants $\lambda_{1},\lambda_{2}>0$ such that $\lambda_{1}+\lambda_{2}<-r/2$ and
$\partial_{yy}H(x,\mu,y)\leq-\lambda_{1},\quad\partial_{xx}H(x,\mu,y)\geq\lambda_{1},\quad|\partial_{xy}H(x,\mu,y)|\leq\lambda_{2}.$ (5.1)
(iii) There exist a constant $\lambda_{3}>0$ such that
$|\partial_{xx}H(x,\mu,y)|\leq\lambda_{3},\quad|\partial_{yy}H(x,\mu,y)|\leq\lambda_{3}.$ (5.2)
(iv) $\partial_{xx}H(x,\mu,y),\partial_{xy}H(x,\mu,y),\partial_{yy}H(x,\mu,y)$ are Lipschitz continuous.
[…]
We introduce the following FBSDE:
$\begin{cases}\mathrm{d}\nabla X^{x,\xi}_{t}=\left[\nabla X^{x,\xi}_{t}\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})+\nabla Y^{x,\xi}_{t}\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right]\mathrm{d}t,\\ \begin{aligned} \mathrm{d}\nabla Y^{x,\xi}_{t}=&-\left[\nabla X^{x,\xi}_{t}\partial_{xx}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})+\nabla Y^{x,\xi}_{t}\partial_{xy}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-r\nabla Y^{x,\xi}_{t}\right]\mathrm{d}t\\ &+\nabla Z^{x,\xi}_{t}\mathrm{d}B_{t},\end{aligned}\\ \nabla X^{x,\xi}_{0}=1.\end{cases}$ (5.3)
Under Assumption 4.1 and 5.1, it’s clear that the above FBSDE satisfies all conditions in Assumption 3.1, so it admits a unique solution in $L_{r}^{2}$. The following theorem tells us that $(\nabla X^{x,\xi},\nabla Y^{x,\xi})$ can be viewed as the derivative of $(X^{x,\xi},Y^{x,\xi})$ with respect to $x$.
Theorem 5.2
For any $x\in\mathbb{R},$ we have
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(X^{x+\delta,\xi}-X^{x,\xi}\right)-\nabla X^{x,\xi}\right\rVert_{r}\to 0,$ (5.4)
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(Y^{x+\delta,\xi}-Y^{x,\xi}\right)-\nabla Y^{x,\xi}\right\rVert_{r}\to 0.$
Proof.  Denote
$\Delta X^{\delta}_{t}=X^{x+\delta,\xi}_{t}-X^{x,\xi}_{t},\quad\Delta Y^{\delta}_{t}=Y^{x+\delta,\xi}_{t}-Y^{x,\xi}_{t},\quad\Delta Z^{\delta}_{t}=Z^{x+\delta,\xi}_{t}-Z^{x,\xi}_{t}.$ (5.5)
By Proposition 4.3, we have
$\left\lVert\left(\Delta X^{\delta},\Delta Y^{\delta},\Delta Z^{\delta}\right)\right\rVert_{r}\leq C\delta,$ (5.6)
where $C$ is a positive constant. Therefore we can define the $L^{2}_{r}$ processes
$\nabla X^{\delta}\triangleq\frac{\Delta X^{\delta}}{\delta},\quad\nabla Y^{\delta}\triangleq\frac{\Delta Y^{\delta}}{\delta},\quad\nabla Z^{\delta}\triangleq\frac{\Delta Z^{\delta}}{\delta}.$ (5.7)
And they satisfy the following FBSDE:
$\begin{cases}\mathrm{d}\nabla X^{\delta}_{t}=\left[\nabla X^{\delta}_{t}\cdot H_{xy}^{\delta}+\nabla Y^{\delta}_{t}\cdot H_{yy}^{\delta}\right]\mathrm{d}t,\\ \begin{aligned} \mathrm{d}\nabla Y^{\delta}_{t}=&-\left[\nabla X^{\delta}_{t}\cdot H_{xx}^{\delta}+\nabla Y^{\delta}_{t}\cdot H_{xy}^{\delta}-r\nabla Y^{\delta}_{t}\right]\mathrm{d}t\\ &+\nabla Z^{\delta}_{t}\mathrm{d}B_{t},\end{aligned}\\ \nabla X^{\delta}_{0}=1,\end{cases}$ (5.8)
where
$H_{xy}^{\delta}=\int_{0}^{1}\partial_{xy}H\left(X_{t}^{x,\xi}+\theta\Delta X^{\delta}_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}+\theta\Delta Y^{\delta}_{t}\right)\mathrm{d}\theta,$ (5.9)
and $H_{xx}^{\delta},H_{yy}^{\delta}$ are defined similarly. Finally, define
$X^{\delta}_{t}\triangleq\nabla X^{\delta}_{t}-\nabla X^{x,\xi}_{t},\quad Y^{\delta}_{t}\triangleq\nabla Y^{\delta}_{t}-\nabla Y^{x,\xi}_{t},\quad Z^{\delta}_{t}\triangleq\nabla Z^{\delta}_{t}-\nabla Z^{x,\xi}_{t}.$ (5.10)
Then we have $(X^{\delta}_{t},Y^{\delta}_{t},Z^{\delta}_{t})$ are uniformly bounded in $L_{r}^{2}$, and they satisfy the following FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{\delta}=\left[H_{xy}^{\delta}X_{t}^{\delta}+H_{yy}^{\delta}Y_{t}^{\delta}+R_{t}^{x}\right]\mathrm{d}t,\\ \begin{aligned} \mathrm{d}Y_{t}^{\delta}&=-\left[H_{xx}^{\delta}X_{t}^{\delta}+H_{xy}^{\delta}Y_{t}^{\delta}-rY_{t}^{\delta}+R_{t}^{y}\right]\mathrm{d}t\\ &\quad+Z_{t}^{\delta}\mathrm{d}B_{t},\end{aligned}\\ X_{0}^{\delta}=0,\end{cases}$ (5.11)
where the remainder terms $R_{t}^{x}$ and $R_{t}^{y}$ are defined as:
$R_{t}^{x}\triangleq$ $\nabla X_{t}^{x,\xi}\left(H_{xy}^{\delta}-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right)$ (5.12)
$+\nabla Y_{t}^{x,\xi}\left(H_{yy}^{\delta}-\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right),$
$R_{t}^{y}\triangleq$ $\nabla X_{t}^{x,\xi}\left(H_{xx}^{\delta}-\partial_{xx}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right)$
$+\nabla Y_{t}^{x,\xi}\left(H_{xy}^{\delta}-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right).$
Applying Itô’s formula to $\mathrm{e}^{-rt}X^{\delta}_{t}Y^{\delta}_{t}$, and noticing that $H_{yy}^{\delta}\leq-\lambda_{1},H_{xx}^{\delta}\geq\lambda_{1},|H_{xy}^{\delta}|\leq\lambda_{2}$, we get that
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left((X_{t}^{\delta})^{2}+(Y_{t}^{\delta})\right)^{2}\mathrm{d}t\right]\leq\frac{2}{r}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\cdot R_{t}^{x}+X_{t}^{\delta}\cdot R_{t}^{y}\right)\mathrm{d}t\right].$ (5.13)
For simplicity, we only estimate the $Y_{t}^{\delta}\cdot\nabla X_{t}^{x,\xi}\left(H_{xx}^{\delta}-\partial_{xx}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right)$ term. Denote
$A_{t}^{\delta}\triangleq\left(H_{xx}^{\delta}-\partial_{xx}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right).$ (5.14)
From the a priori estimate, we know that $Y_{t}^{\delta}$ is uniformly bounded in $L_{r}^{2}$, and $|A_{t}^{\delta}|\leq 2\lambda_{3}$, then there exists a constant $C>0$, such that
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\cdot\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)\mathrm{d}t\right]$ (5.15)
$\leq\left(\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\right)^{2}\mathrm{d}t\right]\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)^{2}\mathrm{d}t\right]\right)^{\frac{1}{2}}$
$\leq C\left(\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)^{2}\mathrm{d}t\right]\right)^{\frac{1}{2}}.$
From the Lipschitz continuity of $\partial_{xx}H$, we know that there exists a constant $\ell>0$, such that
$|A_{t}^{\delta}|\leq\ell\left(|\Delta X^{\delta}_{t}|+|\Delta Y^{\delta}_{t}|\right).$ (5.16)
We consider the following finite measure on $\mathbb{R}_{+}\times\Omega$:
$\mathrm{d}\mathbb{Q}=\mathrm{e}^{-rt}\mathrm{d}t\times\mathrm{d}\mathbb{P}.$ (5.17)
Then we know
$\int|A^{\delta}|^{2}\mathrm{d}\mathbb{Q}\to 0$ (5.18)
as $\delta\to 0$. This shows that under measure $\mathbb{R}$, $|A^{\delta}|^{2}$ converges in measure to 0. Since $(\nabla X^{x,\xi})^{2}$ is integrable under $\mathbb{Q}$, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ converges in measure to 0. Finally, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ is dominated by an integrable function $4\lambda_{3}^{2}(\nabla X^{x,\xi})^{2}$ under measure Q, and by the Dominated Convergence Theorem, we have
$\lim_{\delta\to 0}\int\left(\nabla X^{x,\xi}\cdot A^{\delta}\right)^{2}\mathrm{d}\mathbb{Q}=0.$ (5.19)
This shows that
$\lim_{\delta\to 0}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\cdot\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)\mathrm{d}t\right]=0.$ (5.20)
Now we finish the proof.
Now we suppose $F(x,\mu,y):\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}\rightarrow\mathbb{R}$ satisfies the following conditions:
  • $F$ is of at most quadratic growth and jointly continuous in $(x,\mu,y)$.
  • $\partial_{x}F(x,\mu,y)$ and $\partial_{y}F(x,\mu,y)$ exist and are Lipschitz continuous in $(x,\mu,y)$.
[…]
We can define a function $U$ on $\mathbb{R}\times{\cal P}_{2}$ by
$U(x,\mu)=\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t\right],$ (5.21)
where $\xi\in\mathbb{L}^{2}({\cal F}_{0})$ and ${\cal L}_{\xi}=\mu$. Since we have already established the uniqueness in distribution for the FBSDEs (see [14]), $U$ is well-defined.
We now introduce a lemma about $U$ that will be useful for studying the continuity and differentiability of the value function (1.10).
Lemma 5.3
(i) $\partial_{x}U(x,\mu)$ exists and
[…]
$\qquad+\left.\left.\partial_{y}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\nabla Y^{x,\xi}_{t}\right)\mathrm{d}t\right].$
(ii) $U(x,\mu)$ is jointly continuous in $\mathbb{R}\times{\cal P}_{2}$.
Proof.  (i) We still use the notations in the proof of Theorem 5.2. Fix $x\in\mathbb{R},\xi\in\mathbb{L}^{2}({\cal F}_{0})$ and assume $|\delta|<1$. We have $X_{t}^{x,\xi},{X_{t}^{\xi}},Y_{t}^{x,\xi}$ and $X_{t}^{x+\delta,\xi},Y_{t}^{x+\delta,\xi}$ are all uniformly bounded in $L_{r}^{2}$. From the definition of $U(x,\mu)$, we have
$\frac{U(x+\delta,\mu)-U(x,\mu)}{\delta}=\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(F_{x}^{\delta}\cdot\nabla X^{\delta}_{t}+F_{y}^{\delta}\cdot\nabla Y^{\delta}_{t}\right)\mathrm{d}t\right],$ (5.23)
where
$F_{x}^{\delta}=\int_{0}^{1}\partial_{x}F\left(X_{t}^{x,\xi}+\theta\Delta X^{\delta}_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}+\theta\Delta Y^{\delta}_{t}\right)\mathrm{d}\theta,$ (5.24)
and $F_{y}^{\delta}$ is defined similarly. For simplicity, we let
$\partial_{x}F=\partial_{x}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}),\quad\partial_{y}F=\partial_{y}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}).$ (5.25)
Since $\partial_{x}F(x,\mu,y)$ and $\partial_{y}F(x,\mu,y)$ are Lipschitz continuous, we know that
$\left\lVert F_{x}^{\delta}-F_{x}\right\rVert_{r}\to 0,\quad\left\lVert F_{y}^{\delta}-F_{y}\right\rVert_{r}\to 0$ (5.26)
as $\delta\to 0$. Now we have
$\left\lvert\frac{U(x+\delta,\mu)-U(x,\mu)}{\delta}-\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\partial_{x}F\cdot\nabla X^{x,\xi}_{t}+\partial_{y}F\cdot\nabla Y^{x,\xi}_{t}\right)\mathrm{d}t\right]\right\rvert$ (5.27)
$\leq$ $\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert F_{x}^{\delta}\cdot\nabla X^{\delta}_{t}-\partial_{x}F\cdot\nabla X^{x,\xi}_{t}\right\rvert\mathrm{d}t\right]+\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert F_{y}^{\delta}\cdot\nabla Y^{\delta}_{t}-\partial_{y}F\cdot\nabla Y^{x,\xi}_{t}\right\rvert\mathrm{d}t\right]$
$\triangleq$ $\quad I+II.$
For the first term,
$I\leq$ $\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert F_{x}^{\delta}\cdot\left(\nabla X^{\delta}_{t}-\nabla X^{x,\xi}_{t}\right)\right\rvert\mathrm{d}t\right]+\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert\left(F_{x}^{\delta}-F_{x}\right)\cdot\nabla X^{x,\xi}_{t}\right\rvert\mathrm{d}t\right]$ (5.28)
$\leq$ $\left\lVert F_{x}^{\delta}\right\rVert_{r}\cdot\left\lVert\nabla X^{\delta}_{t}-\nabla X^{x,\xi}_{t}\right\rVert_{r}+\left\lVert F_{x}^{\delta}-F_{x}\right\rVert_{r}\cdot\left\lVert\nabla X^{x,\xi}_{t}\right\rVert_{r}.$
Since $\left\lVert F_{x}^{\delta}\right\rVert_{r}$ and $\left\lVert\nabla X^{x,\xi}_{t}\right\rVert_{r}$ are uniformly bounded, by Theorem 5.2, we know that I tends to 0 as $\delta\rightarrow 0$. The second term can be treated similarly, then we finish the proof.
(ii) Fix $x\in\mathbb{R},\mu\in{\cal P}_{2}$, and consider another $x^{\prime}\in\mathbb{R},\mu^{\prime}\in{\cal P}_{2}$ such that $|x-x^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})<\delta$. And then take $\xi,\xi^{\prime}\in\mathbb{L}^{2}({\cal F}_{0})$, such that ${\cal L}_{\xi}=\mu,{\cal L}_{\xi^{\prime}}=\mu^{\prime}$. Without loss of generality, we can assum $|\delta|<1$ and $|x-x^{\prime}|+(\mathbb{E}[\xi-\xi^{\prime}]^{2})^{1/2}\leq\delta$ , then all processes appeared in the following proof are uniformly bounded in $L_{r}^{2}$ with respect to $\delta$ . We have
$F(X_{t}^{x^{\prime},\xi^{\prime}},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})$ (5.29)
$=$ $F(X_{t}^{x^{\prime},\xi^{\prime}},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})$
$+F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x^{\prime},\xi^{\prime}})$
$+F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})$
$\triangleq$ $\quad I+II+III.$
For the first term,
$F(X_{t}^{x^{\prime},\xi^{\prime}},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})$ (5.30)
$=$ $\int_{0}^{1}\partial_{x}F(X_{t}^{x,\xi}+\theta(X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi}),{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})\mathrm{d}\theta\cdot(X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi})$
$\triangleq$ $A_{t}\cdot(X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi}).$
Since $\partial_{x}F(x,\mu,y)$ is of at most linear growth, we know that $A_{t}$ is uniformly bounded in $L_{r}^{2}$. By Cauchy’s Inequality and Proposition 4.3, We have
$\mathbb{E}\left[\mathrm{e}^{-rt}|I|\mathrm{d}t\right]\leq\left\lVert A_{t}\right\rVert_{r}\cdot\left\lVert X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi}\right\rVert_{r}\leq C_{1}\delta.$ (5.31)
For the second term,
$F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x^{\prime},\xi^{\prime}})$ (5.32)
$=$ $\int_{0}^{1}\tilde{\mathbb{E}}_{{\cal F}_{t}}\left[\partial_{\mu}F\left(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}+\theta(X_{t}^{\xi^{\prime}}-X_{t}^{\xi})},Y_{t}^{x^{\prime},\xi^{\prime}},{\tilde{X}_{t}^{\xi}+\theta(\tilde{X}_{t}^{\xi^{\prime}}-\tilde{X}_{t}^{\xi})}\right)\cdot\left(\tilde{X}_{t}^{\xi^{\prime}}-\tilde{X}_{t}^{\xi}\right)\right]\mathrm{d}\theta.$
[…]
$\mathbb{E}\left[\mathrm{e}^{-rt}|II|\mathrm{d}t\right]\leq C_{2}\delta.$ (5.33)
The approach for III is identical to that for I. Integrating the analyses of I, II, and III, we obtain there exists a constant $C>0$, such that
$|U(x^{\prime},\mu^{\prime})-U(x,\mu)|\leq C\delta.$ (5.34)
This completes the proof.
5.2 Main results
Theorem 5.4
Under Assumption 4.1 and 5.1, the value function satisfies
$\partial_{x}V(x,\mu)=Y_{0}^{x,\xi}.$ (5.35)
Proof.  By the definition of $f(x,\mu,\hat{\alpha}(x,y))$, we have the relationship:
$f(x,\mu,\hat{\alpha}(x,y))=H(x,\mu,y)-\partial_{y}H(x,\mu,y)\cdot y.$ (5.36)
It’s easy to verify that $f(x,\mu,\hat{\alpha}(x,y))$ satisfies all conditions in Lemma 5.3, so $\partial_{x}V(x,\mu)$ exists and
$\partial_{x}V(x,\mu)=\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}$ $\left\{\left[\partial_{x}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\right]\cdot\nabla X^{x,\xi}_{t}\right$. (5.37)
$\left.-\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\cdot\nabla Y^{x,\xi}_{t}\right\}\mathrm{d}t.$
Applying Itô’s formula to $\mathrm{e}^{-rt}\nabla X^{x,\xi}_{t}\cdot Y_{t}^{x,\xi}$, we have
$\mathrm{d}\left(\mathrm{e}^{-rt}\nabla X^{x,\xi}_{t}Y_{t}^{x,\xi}\right)$ $=\mathrm{e}^{-rt}\left[-\nabla X^{x,\xi}_{t}\partial_{x}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})+Y_{t}^{x,\xi}\nabla X^{x,\xi}_{t}\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right$. (5.38)
$\quad\left.+Y_{t}^{x,\xi}\nabla Y^{x,\xi}_{t}\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right]\mathrm{d}t$
$\quad+\mathrm{e}^{-rt}\nabla X^{x,\xi}_{t}Z_{t}^{x}\mathrm{d}B_{t}.$
Note that $\nabla X^{x,\xi}_{0}=1$, and take a sequence $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-rT_{i}}\nabla X^{x,\xi}_{T_{i}}\cdot Y_{T_{i}}^{x,\xi}\right]\to 0.$ (5.39)
Integrating from 0 to $T_{i}$ and taking expectation, after letting $T_{i}\to\infty$ we have:
$Y_{0}^{x,\xi}=\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}$ $\left\{\left[\partial_{x}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\right]\cdot\nabla X^{x,\xi}_{t}\right$. (5.40)
$\left.-\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\cdot\nabla Y^{x,\xi}_{t}\right\}\mathrm{d}t.$
Now we get the desired result.
In preparation for the definition of a viscosity solution of the master equation, we first define the class of test functions used for that purpose.
Definition 5.5
A function $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ is said to be a test function if the quantities:
[…]
and
[…]
are finite, uniformly in $(x,\mu)$ in any compact subset of $\mathbb{R}\times\mathcal{P}_{2}$.
Now, we present the definition of the viscosity solution for master equation (1.5).
Definition 5.6
Suppose that $U\in C(\mathbb{R}\times\mathcal{P}_{2})$ and its partial derivative $\partial_{x}U\in C(\mathbb{R}\times\mathcal{P}_{2})$. Then $U$ is called a viscosity subsolution (resp. supersolution) of PDE (1.5) if, for any $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ and any test function $\Psi$, such that $(x^{0},\mu^{0})$ is a local maximum (resp. minimum) of $U-\Psi$, we have
$\begin{split}rU(x^{0},\mu^{0})\leq&H(x^{0},\mu^{0},\partial_{x}U(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\cdot\partial_{y}H(\tilde{\xi}^{0},\mu^{0},\partial_{x}U(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (5.43)
(respectively
$\begin{split}rU(x^{0},\mu^{0})\geq&H(x^{0},\mu^{0},\partial_{x}U(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\cdot\partial_{y}H(\tilde{\xi}^{0},\mu^{0},\partial_{x}U(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (5.44)
The function $U$ is called a viscosity solution of PDE (1.5) if it is both a viscosity subsolution and a viscosity supersolution.
Theorem 5.7
Under Assumption 4.1 and 5.1, the value function (1.10) is the viscosity solution to the master equation (1.5).
Proof.  We have proved that $V(x,\mu)$ is a continuous function and $\partial_{x}V(x,\mu)=Y_{0}^{x,\xi}$. Since $\partial_{x}V(x,\mu)$ is also a continuous function, by the flow property, we have
$\partial_{x}V(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi},\quad\partial_{x}V(X_{t}^{\xi},{\cal L}_{X_{t}^{\xi}})=Y_{t}^{\xi}.$ (5.45)
We only show that $V$ is a viscosity subsolution of PDE (1.5). A similar argument will show that it is also a viscosity supersolution of (1.5).
Let $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ be a test function and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $V-\Psi$. It’s natural to get that $\partial_{x}V(x^{0},\mu^{0})=\partial_{x}\Psi(x^{0},\mu^{0})$. We assume without loss of generality that $V(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$. And we suppose that
$\begin{split}r\Psi(x^{0},\mu^{0})>&H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\partial_{y}H(\tilde{\xi},\mu^{0},\partial_{x}V(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (5.46)
Notice that
$\begin{split}&H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))\\ =&f(x^{0},\mu^{0},\hat{\alpha}(x^{0},\partial_{x}V(x^{0},\mu^{0})))+\partial_{x}V(x^{0},\mu^{0})\cdot\partial_{y}H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))\\ =&f(x^{0},\mu^{0},\hat{\alpha}(x^{0},\partial_{x}V(x^{0},\mu^{0})))+\partial_{x}\Psi(x^{0},\mu^{0})\cdot\partial_{y}H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0})).\end{split}$ (5.47)
We can get
$\begin{split}r\Psi(x^{0},\mu^{0})>&f(x^{0},\mu^{0},\hat{\alpha}(x^{0},\partial_{x}V(x^{0},\mu^{0})))+\partial_{x}\Psi(x^{0},\mu^{0})\cdot\partial_{y}H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))\\ &+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\partial_{y}H(\tilde{\xi},\mu^{0},\partial_{x}V(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (5.48)
It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$, such that for all $(x,\mu)\in O$,
$\begin{cases}V(x,\mu)\leq\Psi(x,\mu),\\ \begin{aligned} r\Psi(x,\mu)>&f(x,\mu,\hat{\alpha}(x,\partial_{x}V(x,\mu)))+\partial_{x}\Psi(x,\mu)\cdot\partial_{y}H(x,\mu,\partial_{x}V(x,\mu))\\ &+\frac{1}{2}\partial_{xx}\Psi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x,\mu,\tilde{\xi})+\partial_{\mu}\Psi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}V(\tilde{\xi},\mu))\right].\end{aligned}\end{cases}$ (5.49)
Taking an initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ such that ${\cal L}_{\xi^{0}}=\mu^{0}$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (4.1) and (4.2). We denote $\rho_{t}\triangleq\mathcal{L}_{X_{t}^{\xi^{0}}}$. For some $T>0$, let $\tau$ denote the stopping time
$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\rho_{t})\notin O\}\land T.$ (5.50)
By the flow property and dynamic programming principle, we have that
$\Psi(x^{0},\mu^{0})$ $=V(x^{0},\mu^{0})$ (5.51)
$=\mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t$
$\quad+\mathrm{e}^{-r\tau}\int_{\tau}^{\infty}\mathrm{e}^{-r(t-\tau)}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t\bigg]$
$=\mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t+\mathrm{e}^{-r\tau}V(X_{\tau}^{x^{0},\xi^{0}},\rho_{\tau})\bigg]$
$\leq\mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t+\mathrm{e}^{-r\tau}\Psi(X_{\tau}^{x^{0},\xi^{0}},\rho_{\tau})\bigg].$
By the definition of the test function $\Psi$, we can apply Itô’s formula (2.4) to $\mathrm{e}^{-rt}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$, then we get that
$0\leq$ $\ \mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}\bigg(f\Big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}\big(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}}\big)\Big)$ (5.52)
$\quad-r\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})+\partial_{x}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})\cdot\partial_{y}H(X_{t}^{x^{0},\xi^{0}},\rho_{t},Y_{t}^{x^{0},\xi^{0}})$
$\quad+\frac{1}{2}\partial_{xx}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$
$\quad+\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\Big[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})$
$\qquad+\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})\cdot\partial_{y}H(\tilde{X}_{t}^{\xi^{0}},\rho_{t},\tilde{Y}_{t}^{\xi^{0}})\Big]\bigg)\mathrm{d}t\bigg].$
Then by the relationship (5.45), we have that
$0\leq$ $\ \mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}\bigg(-r\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$ (5.53)
$\quad+f\Big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}\big(X_{t}^{x^{0},\xi^{0}},\partial_{x}V(X_{t}^{x^{0},\xi^{0}},\rho_{t})\big)\Big)$
$\quad+\partial_{x}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})\cdot\partial_{y}H\Big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\partial_{x}V(X_{t}^{x^{0},\xi^{0}},\rho_{t})\Big)$
$\quad+\frac{1}{2}\partial_{xx}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$
$\quad+\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\bigg[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})$
$\qquad+\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})\cdot\partial_{y}H\Big(\tilde{X}_{t}^{\xi^{0}},\rho_{t},\partial_{x}V(\tilde{X}_{t}^{\xi^{0}},\rho_{t})\Big)\bigg]\bigg)\mathrm{d}t\bigg],$
which contradicts (5.49). Now we finish the proof.