### v3

# Infinite-Time Mean Field FBSDEs and Viscosity Solutions to Elliptic Master Equations

Abstract
This paper presents a further investigation of the properties of infinite-time mean field FBSDEs and elliptic master equations, which were introduced in [14] as mathematical tools for solving discounted infinite-time mean field games. By establishing the continuous dependence of the FBSDE solutions on their initial values, we prove the flow property of the mean field FBSDEs. And then, we prove that, at the Nash equilibrium, the value function of the representative player constitutes a viscosity solution to the corresponding elliptic master equation. Furthermore, we establish the uniqueness of classical solutions to the elliptic master equation under displacement monotonicity and certain growth assumptions.
Keywords. discounted infinite-time mean field games, infinite-time FBSDEs, elliptic master equations, viscosity solution.
1 Introduction
The study of mean field games was initiated independently by Lasry-Lions [7, 8, 9] and Huang-Malhamé-Caines [6], which is an analysis of limit models for symmetric weakly interacting $(N+1)$-player differential games. The parabolic master equation plays a crucial role in the analysis of mean field games, which was introduced by Lions in lectures [10]. It describes the strategic interaction between a representative player and the collective environment. We refer the reader to [4, 3, 5] for a comprehensive exposition on the subject.
Forward-backward stochastic differential equations (FBSDEs) also serve as a powerful tool for the study of mean field games. The theory of general nonlinear BSDEs was pioneered by Pardoux and Peng [11, 12] in the early 1990s. [13] studied the infinite-time FBSDEs and established connections with quasilinear elliptic PDEs. Recently, [1] extended this framework to the McKean-Vlasov FBSDEs. This paper investigates the properties of infinite-time mean field FBSDEs, and employs them to address the existence of viscosity solutions for elliptic master equations.
In the recent work [14], we proposed the discounted infinite-time mean field games and elliptic master equations, which extends the traditional framework to infinite-time case. In our framework, a representative player interacts with a continuum of other players (also referred to as the population or social equilibrium). Let $\mu_{0}\in{\cal P}(\mathbb{R})$ be the initial state distribution in the society. For the player($x$), a representative player with initial state $x$, the dynamic of its state is given by
$X^{x,\beta}_{t}=x+\int_{0}^{t}b(X^{x,\beta}_{t},\mu_{s},\beta_{s})\mathrm{d}s+B_{t},$ (1.1)
where $\beta$, an $\mathbb{R}$-valued progressively measurable stochastic process, is the strategy of the player($x$) and $(\mu_{t})_{t\geq 0}$ is the population distribution flow start at $\mu_{0}$. Denote by $\beta^{x}$ the optimal strategy of the representative player($x$) derived from the stochastic optimization problem:
$\min_{\beta}J^{\mu}(\beta)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t}^{x,\beta},\mu_{t},\beta_{t})\mathrm{d}t\right].$ (1.2)
Since the population consists of a multitude of individuals, the macroscopic distribution should satisfy
$\mu_{t}(\cdot)=\int\mathbb{P}(X_{t}^{x,\beta^{x}}\in\cdot)\mu_{0}(\mathrm{d}x).$ (1.3)
To solve this fixed-point problem, we introduce the following two systems of FBSDEs. The first one represents the state process of the social equilibrium, while the second denotes the state process of the representative player with initial state $x$.
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\partial_{y}H(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[\partial_{x}{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})-rY_{t}^{\xi}\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (1.4)
$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x.\end{cases}$ (1.5)
Here $r>0$ is the discount factor and
$H(x,\mu,y)\triangleq\min_{a\in\mathbb{R}}\left[b(x,\mu,a)\cdot y+f(x,\mu,a)\right].$ (1.6)
For convenience, we assume that $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a),b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$, and denote by $\hat{\alpha}(x,y)$ the unique minimizer of (1.6). After further assuming that $f$ and $b$ possess good smoothness properties, we can obtain the following relationship:
$\partial_{y}H(x,\mu,y)=b(x,\mu,\hat{\alpha}(x,y)).$ (1.7)
In [14], we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ with ${\cal L}_{\xi}=\mu$ and demonstrate that $\mathcal{V}$ serves as a viscosity solution to the distribution-dependent PDE below:
$\begin{split}r\mathcal{U}(x,\mu)=&\partial_{x}H(x,\mu,\mathcal{U}(x,\mu))+\partial_{y}H(x,\mu,\mathcal{U}(x,\mu))\cdot\partial_{x}\mathcal{U}(x,\mu)+\frac{1}{2}\partial_{xx}\mathcal{U}(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})+\partial_{\mu}\mathcal{U}(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\mathcal{U}(\tilde{\xi},\mu))\right].\end{split}$ (1.8)
Here $\partial_{x},\partial_{xx}$ are standard spatial derivatives, $\partial_{\mu},\partial_{\tilde{x}\mu}$ are ${\cal W}_{2}$-Wasserstein derivatives, $\tilde{\xi}$ is a random variable with law $\mu$ and $\tilde{\mathbb{E}}$ is the expectation with respect to its law. We also introduce the elliptic master equation:
$\begin{split}rU(x,\mu)=&H(x,\mu,\partial_{x}U(x,\mu))+\frac{1}{2}\partial_{xx}U(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})+\partial_{\mu}U(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}U(\tilde{\xi},\mu))\right].\end{split}$ (1.9)
We have proved that if this master equation admits a classical solution with sufficient regularity, we can derive the following representation for FBSDEs (1.4) and (1.5):
$Y_{t}^{\xi}=\partial_{x}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{\xi}=\partial_{xx}U(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}}),$ (1.10)
$Y_{t}^{x,\xi}=\partial_{x}U(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}}),\quad Z_{t}^{x,\xi}=\partial_{xx}U(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}}).$ (1.11)
In this paper, we make above relation rigorous and proceed to study the properties of the elliptic master equation. We prove that solutions to FBSDEs (1.4) and (1.5) possess the flow property:
$X_{t}^{x,\xi}|_{x=\xi}=X_{t}^{\xi},\quad Y_{t}^{x,\xi}|_{x=\xi}=Y_{t}^{\xi},\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$},$ (1.12)
and
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (1.13)
Furthermore, we prove that the value function of the representative player
$V(x,\mu)=\mathbb{E}\bigg[\int_{0}^{+\infty}\mathrm{e}^{-rt}f\big(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},\hat{\alpha}(X_{t}^{x,\xi},Y_{t}^{x,\xi})\big)\mathrm{d}t\bigg]$ (1.14)
is a viscosity solution to elliptic master equation (1.9), and satisfies the relationship
$\partial_{x}V(x,\mu)=Y_{0}^{x,\xi}.$ (1.15)
Finally, we address the uniqueness of classical solutions to the elliptic master equation. We introduce a displacement monotonicity condition (see [5]), which states that for any square integrable random variables $\xi_{1},\xi_{2}$, the following inequality holds:
$\mathbb{E}\left[(\xi_{1}-\xi_{2})\left(\partial_{x}U(\xi_{1},\mathcal{L}_{\xi_{1}})-\partial_{x}U(\xi_{2},\mathcal{L}_{\xi_{2}})\right)\right]\geq 0.$ (1.16)
Under suitable growth conditions and this displacement monotonicity condition, we establish the uniqueness of classical solutions to the elliptic master equation.
2 Preliminaries
We will use a filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ endowed with a Brownian motion $B$. Its filtration $\mathbb{F}\triangleq(\mathcal{F}_{t})_{t\geq 0}$ is augmented by all $\mathbb{P}$-null sets and a sufficiently rich sub-$\sigma$-algebra $\mathcal{F}_{0}$ independent of $B$, such that it can support any measure on $\mathbb{R}$ with finite second moment.
Let $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ be a copy of the filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with corresponding Brownian motion $B^{\prime}$, define the larger filtered probability space by
$\tilde{\Omega}\triangleq\Omega\times\Omega^{\prime},\quad\tilde{\mathcal{F}}\triangleq\mathcal{F}\otimes\mathcal{F}^{\prime}\quad\tilde{\mathbb{F}}=\{\tilde{\cal F}_{t}\}_{t\geq 0}\triangleq\{{\cal F}_{t}\otimes{\cal F}^{\prime}_{t}\}_{t\geq 0},\quad\tilde{\mathbb{P}}\triangleq\mathbb{P}\otimes\mathbb{P}^{\prime},\quad\tilde{\mathbb{E}}\triangleq\mathbb{E}^{\tilde{\mathbb{P}}}$. (2.1)
[…]
Let ${\cal P}\triangleq{\cal P}(\mathbb{R})$ be the set of all probability measures on $\mathbb{R}$ and let ${\cal P}_{p}(p\geq 1)$ denote the set of $\mu\in{\cal P}$ with finite $p$-th moment. For any sub-$\sigma$-field $\cal G$$\subset{\cal F}$ and $\mu\in{\cal P}_{p}$, we define $\mathbb{L}^{p}({\cal G})$ to be the set of $\mathbb{R}$-valued, $\cal G$}-measurable, and $p$-integrable random variables $\xi$ , and $\mathbb{L}^{p}({\cal G};\mu)$ to be the set of $\xi\in\mathbb{L}^{p}({\cal G})$ such that the law ${\cal L}_{\xi}=\mu$ . For any $\mu,\nu\in{\cal P}_{p}$, we define the ${\cal W}_{p}$–Wasserstein distance between them as follows:
${\cal W}_{p}(\mu,\nu)\triangleq\inf\Big\{\big(\mathbb{E}[|\xi-\eta|^{q}]\big)^{1/q}:\mbox{for all $\xi\in\mathbb{L}^{p}({\cal F};\mu)$, $\eta\in\mathbb{L}^{p}({\cal F};\nu)$}\Big\}.$
Due to our interest in discounted infinite-time mean field games, for any $K\in\mathbb{R}$, we denote by $L^{2}_{K}(t_{0},\infty,\mathbb{R})$ the Hilbert space of all $\mathbb{R}$-valued adapted stochastic process ($v_{t}$) start from $t_{0}$ such that
$\mathbb{E}\left[\int_{t_{0}}^{\infty}e^{-Kt}|v_{t}|^{2}dt\right]<+\infty.$ (2.2)
To simplify, we set $L^{2}_{K}\triangleq L^{2}_{K}(0,\infty,\mathbb{R})$ and define the exponentially weighted $L^{2}$ norm
$\|v\|_{K}^{2}\triangleq\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-Kt}|v_{t}|^{2}\mathrm{d}t\right].$ (2.3)
For each $\mathcal{F}_{0}$-measurable square integrable random variable $\xi$ , we consider the following infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.4)
Here $G,F:\mathbb{R}^{+}\times\mathbb{R}^{2}\times{\cal P}_{2}\to\mathbb{R}$ are two measurable functions, and satisfy the following assumptions:
Assumption 2.1
(i) There exists a positive constant $\ell$ such that for any $x,x^{\prime},y,y^{\prime}\in\mathbb{R},\mu,\mu^{\prime}\in{\cal P}_{2}$
$|G(t,x,y,\mu)$ $-G(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.5)
$\leq\ell(|x-x^{\prime}|+|y-y^{\prime}|+\mathcal{W}_{2}(\mu,\mu^{\prime})).\quad\mathrm{a.s.}$
(ii) There exist constants $0<K<2\kappa$ such that for any $t\geq 0$ and any square integrable random variables $X,X^{\prime},Y,Y^{\prime}$
$\mathbb{E}\left[-K\hat{X}\hat{Y}-\hat{X}(F(t,U)-F(t,U^{\prime}))+\hat{Y}(B(t,U)-B(t,U^{\prime}))\right]$ (2.6)
$\leq-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}}).$
The following lemma states the existence and uniqueness of a solution to FBSDE(2.4). For the detailed proof, we refer the reader to ([1], Theorem 2.1).
Lemma 2.2
Under Assumption 2.1, the FBSDE (2.4) admits a unique solution in $L_{K}^{2}$.
We introduce the Wasserstein space and differential calculus on Wasserstein space. For a ${\cal W}_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$, its ${\cal W}_{2}$-Wasserstein derivatives[4](also called Lions-derivative), takes the form $\partial_{\mu}U:(\mu,\tilde{x})\in{\cal P}_{2}\times\mathbb{R}\to\mathbb{R}$ and satisfies:
$U({\cal L}_{\xi+\eta})-U(\mu)=\mathbb{E}\big[\langle\partial_{\mu}U(\mu,\xi),\eta\rangle\big]+o(\|\eta\|_{2}),\ \forall\ \xi\in\mathbb{L}^{2}(\mathcal{F};\mu),\eta\in\mathbb{L}^{2}(\mathcal{F}).$ (2.7)
Let $C^{0}({\cal P}_{2})$ denote the set of ${\cal W}_{2}$-continuous functions $U:{\cal P}_{2}\to\mathbb{R}$. For $C^{1}({\cal P}_{2})$, we mean the space of functions $U\in C^{0}({\cal P}_{2})$ such that $\partial_{\mu}U$ exists and is continuous on ${\cal P}_{2}\times\mathbb{R}$, which is uniquely determined by (2.7). Let $C^{2,1}(\mathbb{R}\times{\cal P}_{2})$ denote the set of continuous functions $U:\mathbb{R}\times{\cal P}_{2}\to\mathbb{R}$ such that $\partial_{x}U,\partial_{xx}U$ exist and are joint continuous on $\mathbb{R}\times{\cal P}_{2}$, $\partial_{\mu}U,\partial_{x\mu}U,\partial_{\tilde{x}\mu}U$ exist and are continuous on $\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}$.
Finally, we consider the space $\Theta\triangleq[0,T]\times\mathbb{R}\times{\cal P}_{2}$ for some $T>0$, and let $C^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in C^{1,2,1}(\Theta)$ is the Itô’s formula[2, 4]. For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,T]\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,T]\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$. Suppose that for every compact subset ${\cal K}\subset\mathbb{R}\times{\cal P}_{2}$, it holds that:
[…]
We have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big[\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\partial_{xx}U(\sigma_{t}^{1})^{2}\Big](t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big(\tilde{\mathbb{E}}_{{\cal F}_{t}}\big[\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{\sigma}_{t}^{2})^{2}\big]\Big)\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.9)
Here $\tilde{\mathbb{E}}_{{\cal F}_{t}}$ is the conditional expectations given ${\cal F}_{t}$ corresponding to the probability measure $\tilde{\mathbb{P}}$ and the process $(\tilde{X}^{2}_{t},\tilde{b}^{2}_{t},\tilde{\sigma}_{t}^{2})_{0\leq t\leq T}$ is a copy of the process $(X^{2}_{t},b^{2}_{t},\sigma_{t}^{2})_{0\leq t\leq T}$, on a copy $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime},\mathbb{F}^{\prime})$ of the probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$.
3 Properties of mean field FBSDEs
In this section, for some $\xi\in\mathbb{L}^{2}({\cal F}_{0})$, we investigate the following mean field FBSDEs:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\partial_{y}H(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[\partial_{x}{H}(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi})-rY_{t}^{\xi}\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi,\end{cases}$ (3.1)
$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\partial_{y}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x,\xi}=x.\end{cases}$ (3.2)
In [14], we provided sufficient conditions on $b(x,\mu,a)$ and $f(x,\mu,a)$ to ensure the above two FBSDEs admit unique solutions. To obtain further properties of these solutions , we require the following assumptions on $H$:
Assumption 3.1
(i) $H(x,\mu,y)$ is jointly continuous and and all second-order partial derivatives exist.
(ii) $\partial_{y}H(x,\mu,y)$ and $\partial_{x}H(x,\mu,y)$ are Lipschitz continuous in $(x,\mu,y)$. More specifically, there exists a constant $\ell>0$ such that
$|\partial_{y}H(x,\mu,y)-\partial_{y}H(x^{\prime},\mu^{\prime},y^{\prime})|\leq\ell\left(|x-x^{\prime}|+|y-y^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})\right),$ (3.3)
$|\partial_{x}H(x,\mu,y)-\partial_{x}H(x^{\prime},\mu^{\prime},y^{\prime})|\leq\ell\left(|x-x^{\prime}|+|y-y^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})\right).$
(iii)There exist constants $\kappa,C_{0}>0$ , such that $C_{0}+r/2<\kappa$ and
$-(x-x^{\prime})\left[\partial_{x}H(x,\mu,y)-\partial_{x}H(x^{\prime},\mu^{\prime},y^{\prime})\right]+(y-y^{\prime})\left[\partial_{y}H(x,\mu,y)-\partial_{y}H(x^{\prime},\mu^{\prime},y^{\prime})\right]$ (3.4)
$\leq-\kappa\left(|x-x^{\prime}|^{2}+|y-y^{\prime}|^{2}\right)+C_{0}{\cal W}_{2}^{2}(\mu,\mu^{\prime}).$
Theorem 3.2
Under Assumption 3.1, both FBSDE (3.1) and (3.2) have unique solutions in $L_{r}^{2}$.
Proof.  For FBSDE (3.1), it’s clear that the conditions in Assumption 2.1 (i) and (ii) are satisfied. Taking four arbitrary square integrable random variables $X,X^{\prime},Y,Y^{\prime}$, we have
$\mathbb{E}\biggl[$ $-r\hat{X}\cdot\hat{Y}$ (3.5)
$-\hat{X}\left[\partial_{x}H(X,{\cal L}_{X},Y)-\partial_{x}H(X^{\prime},{\cal L}_{X^{\prime}},Y^{\prime})-r\hat{Y}\right$]
$+\hat{Y}\left[\partial_{y}H(X,{\cal L}_{X},Y)-\partial_{y}H(X^{\prime},{\cal L}_{X^{\prime}},Y^{\prime})\right]\biggr]$
$\leq$ $-\kappa\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right]+C_{0}\mathcal{W}_{2}^{2}({\cal L}_{X},{\cal L}_{X^{\prime}})$
$\leq$ $-(\kappa-C_{0})\mathbb{E}\left[\hat{X}^{2}+\hat{Y}^{2}\right],$
where $\hat{X}\triangleq X-X^{\prime},\hat{Y}\triangleq Y-Y^{\prime}$. Since $\kappa-C_{0}>r/2$, the conditions in Assumption 2.1 (iii) is satisfied. Then, FBSDE (3.1) has a unique solution in $L_{r}^{2}$.
After solving FBSDE (3.1), we substitute its solution ${\cal L}_{X_{t}^{\xi}}$ into FBSDE (3.2), and it can similarly be shown that there exists a unique solution to FBSDE (3.1).
The following proposition informs us that the solution of the mean field FBSDE exhibits favorable continuous dependence on the initial value, which is of paramount importance for our subsequent research.
Proposition 3.3
For FBSDEs (3.1) and (3.2), assuming all conditions in Assumption 3.1 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\xi,\xi^{\prime}\in\mathbb{L}^{2}({\cal F}_{0})$, there exists a constant $C>0$, such that
$\mathbb{E}\left[|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}|^{2}\right]+\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\leq C\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]$ (3.6)
and
$|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}|^{2}+\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\leq C\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (3.7)
Proof.  Set
$\begin{split}&\hat{X}^{\xi}=X^{\xi}-X^{\xi^{\prime}},\quad\hat{Y}^{\xi}=Y^{\xi}-Y^{\xi^{\prime}},\quad\hat{Z}^{\xi}=Z^{\xi}-Z^{\xi^{\prime}},\\ &\hat{X}^{x,\xi}=X^{x,\xi}-X^{x^{\prime},\xi^{\prime}},\quad\hat{Y}^{x,\xi}=Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}},\quad\hat{Z}^{x,\xi}=Z^{x,\xi}-Z^{x^{\prime},\xi^{\prime}}.\end{split}$ (3.8)
$C_{1},C_{2},C_{3},C_{4},C_{5},C_{6},C_{7}$ appeared in the following proof are all positive constants.
Applying Itô’s formula to $\mathrm{e}^{-rt}|Y^{\xi}_{t}-Y^{\xi^{\prime}}_{t}|^{2}$, and taking a sequence of $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-rt}|Y^{\xi}_{T_{i}}-Y^{\xi^{\prime}}_{T_{i}}|^{2}\right]\rightarrow 0,$ (3.9)
we get that
$\mathbb{E}\left[\left|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}\right|^{2}\right]=$ $\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[r\left|Y_{t}^{\xi}-Y_{t}^{\xi^{\prime}}\right|^{2}$ (3.10)
$\quad+2\hat{Y}^{\xi}\cdot\left(\partial_{x}H\left(X_{t}^{\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{\xi}\right)-\partial_{x}H\left(X_{t}^{\xi^{\prime}},\mathcal{L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{\xi^{\prime}}\right)-r\hat{Y}^{\xi}\right)$
$\quad-\left|Z_{t}^{\xi}-Z_{t}^{\xi^{\prime}}\right|^{2}\biggr]\mathrm{d}t$
$\leq$ $C_{1}\left(\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\right).$
Applying Itô’s formula to $\mathrm{e}^{-rt}\hat{X}^{\xi}\hat{Y}^{\xi}$, we get
$-\mathbb{E}\left[\hat{X}^{\xi}_{0}\hat{Y}^{\xi}_{0}\right]=$ $\ \mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[-r\hat{X}^{\xi}_{t}\hat{Y}^{\xi}_{t}$ (3.11)
$+\hat{Y}^{\xi}_{t}\left(\partial_{y}H(X^{\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{\xi}_{t})-\partial_{y}H(X^{\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{\xi^{\prime}}_{t})\right)$
$-\hat{X}^{\xi}_{t}\left(\partial_{x}H(X^{\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{\xi}_{t})-\partial_{x}H(X^{\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{\xi^{\prime}}_{t})-r\hat{Y}^{\xi}_{t}\right)\biggr]\mathrm{d}t$
$\leq$ $-\frac{r}{2}\left(\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\right).$
Then, we have
$\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\leq$ $\frac{2}{r}\mathbb{E}\left[\hat{X}^{\xi}_{0}\hat{Y}^{\xi}_{0}\right]$ (3.12)
$\leq$ $\frac{1}{2C_{1}}\mathbb{E}\left[\left|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}\right|^{2}\right]+\frac{32C_{1}}{r^{2}}\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right].$
Combining (3.10) and (3.12), we get that
$\mathbb{E}\left[\left|Y_{0}^{\xi}-Y_{0}^{\xi^{\prime}}\right|^{2}\right]\leq C_{2}\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right].$ (3.13)
Substituting this into (3.12), we obtain
$\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{\xi}-Y^{\xi^{\prime}}\right\rVert_{r}^{2}\leq C_{3}\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right].$ (3.14)
Next, we employ the same approach to FBSDE (3.2). Applying Itô’s formula to $\mathrm{e}^{-rt}|Y^{x,\xi}_{t}-Y^{x^{\prime},\xi^{\prime}}_{t}|^{2}$, we get that
$\mathbb{E}\left[\left|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}\right|^{2}\right]=$ $\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[r\left|Y_{t}^{x,\xi}-Y_{t}^{x^{\prime},\xi^{\prime}}\right|^{2}$ (3.15)
$\quad+2\hat{Y}^{x,\xi}\cdot\left(\partial_{x}H\left(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}\right)-\partial_{x}H\left(X_{t}^{x^{\prime},\xi^{\prime}},\mathcal{L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}}\right)-r\hat{Y}^{x,\xi}\right)$
$\quad-\left|Z_{t}^{x,\xi}-Z_{t}^{x^{\prime},\xi^{\prime}}\right|^{2}\biggr]\mathrm{d}t$
$\leq$ $C_{4}\left(\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}\right).$
Applying Itô’s formula to $\mathrm{e}^{-rt}\hat{X}^{x,\xi}\hat{Y}^{x,\xi}$, we can get that
$-\mathbb{E}\left[\hat{X}^{x,\xi}_{0}\hat{Y}^{x,\xi}_{0}\right]=$ $\ \mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\biggl[-r\hat{X}^{x,\xi}_{t}\hat{Y}^{x,\xi}_{t}$ (3.16)
$+\hat{Y}^{x,\xi}_{t}\left(\partial_{y}H(X^{x,\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{x,\xi}_{t})-\partial_{y}H(X^{x^{\prime},\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{x^{\prime},\xi^{\prime}}_{t})\right)$
$-\hat{X}^{x,\xi}_{t}\left(\partial_{x}H(X^{x,\xi}_{t},\mathcal{L}_{X^{\xi}_{t}},Y^{x,\xi}_{t})-\partial_{x}H(X^{x^{\prime},\xi^{\prime}}_{t},\mathcal{L}_{X^{\xi^{\prime}}_{t}},Y^{x^{\prime},\xi^{\prime}}_{t})-r\hat{Y}^{x,\xi}_{t}\right)\biggr]\mathrm{d}t$
$\leq$ $-\frac{r}{2}\left(\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\right)+C_{0}\left\lVert X^{\xi}-X^{\xi^{\prime}}\right\rVert_{r}^{2}.$
Then, we have
$\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\leq$ $\frac{1}{2C_{4}}\mathbb{E}\left[\left|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}\right|^{2}\right]+C_{5}\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (3.17)
Substituting it into (3.15), we have
$\mathbb{E}\left[\left|Y_{0}^{x,\xi}-Y_{0}^{x^{\prime},\xi^{\prime}}\right|^{2}\right]\leq$ $C_{6}\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (3.18)
At last we come back to (3.17), we get that
$\left\lVert X^{x,\xi}-X^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}+\left\lVert Y^{x,\xi}-Y^{x^{\prime},\xi^{\prime}}\right\rVert_{r}^{2}\leq$ $C_{7}\left(|x-x^{\prime}|^{2}+\mathbb{E}\left[|{\xi}-{\xi^{\prime}}|^{2}\right]\right).$ (3.19)
Now, we complete the proof.
Benefiting from the continuous dependence on initial values of the solutions to the mean field FBSDEs, we can demonstrate the flow property. This precisely reflects the essence of the mean field game: the aggregation of innumerable representative players constitutes the equilibrium state of the system.
Theorem 3.4
Let Assumption 3.1 hold. For any $\xi\in\mathbb{L}^{2}({\cal F}_{0})$, we have
$X_{t}^{x,\xi}|_{x=\xi}=X_{t}^{\xi},\quad Y_{t}^{x,\xi}|_{x=\xi}=Y_{t}^{\xi},\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$},$ (3.20)
and
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (3.21)
Proof.  We prove this theorem in two steps.
Step 1. We first assume $\xi$ is discrete, that is
$\xi=\sum_{i=1}^{n}x_{i}I_{A_{i}},$ (3.22)
where $x_{i}\in\mathbb{R}$ are constants and $\{A_{i}\}\in{\cal F}_{0}$ is a partition of $\Omega$. For each $i\in\{1,2,\cdots,n\}$, we can solve the following FBSDE in $L_{r}^{2}$:
$\begin{cases}\mathrm{d}X_{t}^{x_{i},\xi}=\partial_{y}H(X_{t}^{x_{i},\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x_{i},\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x_{i},\xi}=-\left[\partial_{x}{H}(X_{t}^{x_{i},\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x_{i},\xi})-rY_{t}^{x_{i},\xi}\right]\mathrm{d}t+Z_{t}^{x_{i}}\mathrm{d}B_{t},\\ X_{0}^{x_{i}}=x_{i}.\end{cases}$ (3.23)
Now, we define
$X_{t}(\omega)\triangleq\sum_{i=1}^{n}X_{t}^{x_{i},\xi}(\omega)\cdot I_{A_{i}}(\omega),$ (3.24)
and similarly define $Y$ and $Z$ in the same manner. Then, we know $(X,Y,Z)\in L_{r}^{2}$.
Multiplying both sides of the individual integral equations by $I_{A_{i}}$ and summing over $i$, we derive the aggregated equations. For the forward process:
$\begin{split}X_{t}&=\sum_{i=1}^{n}X_{t}^{x_{i},\xi}I_{A_{i}}\\ &=\sum_{i=1}^{n}\left(x_{i}+\int_{0}^{t}\partial_{y}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})\mathrm{d}s+B_{t}\right)I_{A_{i}}\\ &=\sum_{i=1}^{n}x_{i}I_{A_{i}}+\int_{0}^{t}\sum_{i=1}^{n}\partial_{y}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})I_{A_{i}}\mathrm{d}s+\sum_{i=1}^{n}I_{A_{i}}B_{t}\\ &=\xi+\int_{0}^{t}\partial_{y}H(X_{s},\mathcal{L}_{X_{s}^{\xi}},Y_{s})\mathrm{d}s+B_{t}.\end{split}$ (3.25)
For the backward process:
$\begin{split}Y_{t}&=\sum_{i=1}^{n}Y_{t}^{x_{i},\xi}I_{A_{i}}\\ &=\sum_{i=1}^{n}\left(Y_{T}^{x_{i},\xi}+\int_{t}^{T}\left[\partial_{x}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})-rY_{s}^{x_{i},\xi}\right]\mathrm{d}s-\int_{t}^{T}Z_{s}^{x_{i},\xi}\mathrm{d}B_{s}\right)I_{A_{i}}\\ &=\sum_{i=1}^{n}Y_{T}^{x_{i},\xi}I_{A_{i}}+\int_{t}^{T}\sum_{i=1}^{n}\left[\partial_{x}H(X_{s}^{x_{i},\xi},\mathcal{L}_{X_{s}^{\xi}},Y_{s}^{x_{i},\xi})-rY_{s}^{x_{i},\xi}\right]I_{A_{i}}\mathrm{d}s-\int_{t}^{T}\sum_{i=1}^{n}Z_{s}^{x_{i},\xi}I_{A_{i}}\mathrm{d}B_{s}\\ &=Y_{T}+\int_{t}^{T}\left[\partial_{x}H(X_{s},\mathcal{L}_{X_{s}^{\xi}},Y_{s})-rY_{s}\right]\mathrm{d}s-\int_{t}^{T}Z_{s}\mathrm{d}B_{s}.\end{split}$ (3.26)
Thus, the aggregated processes $(X,Y,Z)$ satisfy the following FBSDE:
$\begin{cases}\mathrm{d}X_{t}=\partial_{y}H(X_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-\left[\partial_{x}{H}(X_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t})-rY_{t}\right]\mathrm{d}t+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (3.27)
We have proved that, given $\mathcal{L}_{X_{t}^{\xi}}$, the above FBSDE has a unique solution in $L_{r}^{2}$. So
$\left\lVert X-X^{\xi}\right\rVert_{r}=\left\lVert Y-Y^{\xi}\right\rVert_{r}=0$ (3.28)
and
$\mathbb{E}\left[|Y_{0}-Y_{0}^{\xi}|^{2}\right]=0.$ (3.29)
Then, we have
$X_{t}^{x,\xi}(\omega)|_{x=\xi(\omega)}=X_{t}^{\xi}(\omega),\quad Y_{t}^{x,\xi}(\omega)|_{x=\xi(\omega)}=Y_{t}^{\xi}(\omega),\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$},$ (3.30)
and
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (3.31)
Step 2. In the general case, let $\xi_{n}\in\mathbb{L}^{2}({\cal F}_{0})$ be a discrete approximation of $\xi$ such that $\mathbb{E}\left[\left\lvert\xi_{n}-\xi\right\rvert^{2}\right]\rightarrow 0$. By Step 1, we have
$X_{t}^{x,\xi_{n}}(\omega)|_{x=\xi_{n}(\omega)}=X_{t}^{\xi_{n}}(\omega),\quad Y_{t}^{x,\xi_{n}}(\omega)|_{x=\xi_{n}(\omega)}=Y_{t}^{\xi_{n}}(\omega)\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$}.$ (3.32)
Applying Proposition 3.3, we have
$\left\lVert X^{\xi_{n}}-X^{\xi}\right\rVert_{r}\to 0,\quad\left\lVert Y^{\xi_{n}}-Y^{\xi}\right\rVert_{r}\to 0.$ (3.33)
On the other hand,
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(X_{t}^{\xi(\omega),\xi}-X_{t}^{\xi_{n}(\omega),\xi_{n}}\right)^{2}\mathrm{d}t\right]\leq$ $\mathbb{E}\left[\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(X_{t}^{\xi(\omega),\xi}-X_{t}^{\xi_{n}(\omega),\xi_{n}}\right)^{2}\mathrm{d}t|{\cal F}_{0}\right]\right]$ (3.34)
$\leq$ $C\mathbb{E}\left[|\xi(\omega)-\xi_{n}(\omega)|^{2}+\mathbb{E}\left[|\xi-\xi_{n}|^{2}\right]\right]\rightarrow 0.$
So we have $X_{t}^{\xi}$ and $X_{t}^{x,\xi}|_{x=\xi}$ are the same in $L_{r}^{2}$, The same holds for $Y_{t}^{\xi}$ and $Y_{t}^{x,\xi}|_{x=\xi}$. Then we conclude that
$X_{t}^{x,\xi}|_{x=\xi}=X_{t}^{\xi},\quad Y_{t}^{x,\xi}|_{x=\xi}=Y_{t}^{\xi},\quad~\mbox{for $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.e. $(t,\omega)$}.$ (3.35)
In addition, by Proposition 3.3, we also have
$Y_{0}^{x,\xi_{n}}|_{x=\xi_{n}}=Y_{0}^{\xi_{n}},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (3.36)
and
$\mathbb{E}\left[|Y_{0}^{\xi}-Y_{0}^{\xi_{n}}|^{2}\right]\leq C\mathbb{E}\left[|\xi-\xi_{n}|^{2}\right]\to 0,$ (3.37)
$\mathbb{E}\left[|Y_{0}^{\xi(\omega),\xi}-Y_{0}^{\xi_{n}(\omega),\xi_{n}}|^{2}\right]=\mathbb{E}\left[\mathbb{E}\left[|Y_{0}^{\xi(\omega),\xi}-Y_{0}^{\xi_{n}(\omega),\xi_{n}}|^{2}|{{\cal F}_{0}}\right]\right]$
$\leq C\mathbb{E}\left[|\xi(\omega)-\xi_{n}(\omega)|^{2}+\mathbb{E}\left[|\xi-\xi_{n}|^{2}\right]\right]\rightarrow 0.$
Thus, we have
$Y_{0}^{x,\xi}|_{x=\xi}=Y_{0}^{\xi},\quad~\mbox{for $\mathrm{d}\mathbb{P}$-a.e. $\omega$}.$ (3.38)
4 Solutions to elliptic master equations
In this section, we will first prove the main result of this paper: the value function (1.14) is a viscosity solution to the elliptic master equation (1.9). And then we show the uniqueness of classical solutions to the elliptic master equation under suitable growth conditions and the displacement monotonicity condition.
4.1 Differentiability of mean field FBSDEs
In order to prove the main theorem, we first establish the differentiability of the solutions to the mean field FBSDEs. For this purpose, we require the following assumptions:
Assumption 4.1
(i) $H(x,\mu,y)$ has at most quadratic growth.
(ii) There exist a constant $\lambda_{1}>r/2$ such that
$\partial_{yy}H(x,\mu,y)\leq-\lambda_{1},\quad\partial_{xx}H(x,\mu,y)\geq\lambda_{1}.$ (4.1)
(iii) There exist a constant $\lambda_{2}>0$ such that
$|\partial_{xx}H(x,\mu,y)|\leq\lambda_{2},\quad|\partial_{yy}H(x,\mu,y)|\leq\lambda_{2},\quad|\partial_{xy}H(x,\mu,y)|\leq\lambda_{2}.$ (4.2)
(iv) $\partial_{xx}H(x,\mu,y),\partial_{xy}H(x,\mu,y),\partial_{yy}H(x,\mu,y)$ are Lipschitz continuous.
[…]
We introduce the following FBSDE:
$\begin{cases}\mathrm{d}\nabla X^{x,\xi}_{t}=\left[\nabla X^{x,\xi}_{t}\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})+\nabla Y^{x,\xi}_{t}\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right]\mathrm{d}t,\\ \begin{aligned} \mathrm{d}\nabla Y^{x,\xi}_{t}=&-\left[\nabla X^{x,\xi}_{t}\partial_{xx}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})+\nabla Y^{x,\xi}_{t}\partial_{xy}{H}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-r\nabla Y^{x,\xi}_{t}\right]\mathrm{d}t\\ &+\nabla Z^{x,\xi}_{t}\mathrm{d}B_{t},\end{aligned}\\ \nabla X^{x,\xi}_{0}=1.\end{cases}$ (4.3)
Under Assumption 3.1 and 4.1, it’s clear that the above FBSDE satisfies all conditions in Assumption 2.1, so it admits a unique solution in $L_{r}^{2}$. The following theorem tells us that $(\nabla X^{x,\xi},\nabla Y^{x,\xi})$ can be viewed as the derivative of $(X^{x,\xi},Y^{x,\xi})$ with respect to $x$.
Theorem 4.2
For any $x\in\mathbb{R}$, we have
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(X^{x+\delta,\xi}-X^{x,\xi}\right)-\nabla X^{x,\xi}\right\rVert_{r}=0,$ (4.4)
$\lim_{\delta\to 0}\left\lVert\frac{1}{\delta}\left(Y^{x+\delta,\xi}-Y^{x,\xi}\right)-\nabla Y^{x,\xi}\right\rVert_{r}=0.$
Proof.  Denote
$\Delta X^{\delta}_{t}=X^{x+\delta,\xi}_{t}-X^{x,\xi}_{t},\quad\Delta Y^{\delta}_{t}=Y^{x+\delta,\xi}_{t}-Y^{x,\xi}_{t},\quad\Delta Z^{\delta}_{t}=Z^{x+\delta,\xi}_{t}-Z^{x,\xi}_{t}.$ (4.5)
By Proposition 3.3, we have
$\left\lVert\left(\Delta X^{\delta},\Delta Y^{\delta},\Delta Z^{\delta}\right)\right\rVert_{r}\leq C\delta,$ (4.6)
where $C$ is a positive constant. Therefore we can define the $L^{2}_{r}$ processes
$\nabla X^{\delta}\triangleq\frac{\Delta X^{\delta}}{\delta},\quad\nabla Y^{\delta}\triangleq\frac{\Delta Y^{\delta}}{\delta},\quad\nabla Z^{\delta}\triangleq\frac{\Delta Z^{\delta}}{\delta},$ (4.7)
and they satisfy the following FBSDE:
$\begin{cases}\mathrm{d}\nabla X^{\delta}_{t}=\left[\nabla X^{\delta}_{t}\cdot H_{xy}^{\delta}+\nabla Y^{\delta}_{t}\cdot H_{yy}^{\delta}\right]\mathrm{d}t,\\ \begin{aligned} \mathrm{d}\nabla Y^{\delta}_{t}=&-\left[\nabla X^{\delta}_{t}\cdot H_{xx}^{\delta}+\nabla Y^{\delta}_{t}\cdot H_{xy}^{\delta}-r\nabla Y^{\delta}_{t}\right]\mathrm{d}t\\ &+\nabla Z^{\delta}_{t}\mathrm{d}B_{t},\end{aligned}\\ \nabla X^{\delta}_{0}=1,\end{cases}$ (4.8)
where
$H_{xy}^{\delta}=\int_{0}^{1}\partial_{xy}H\left(X_{t}^{x,\xi}+\theta\Delta X^{\delta}_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}+\theta\Delta Y^{\delta}_{t}\right)\mathrm{d}\theta,$ (4.9)
and $H_{xx}^{\delta},H_{yy}^{\delta}$ are defined similarly. Finally, define
$X^{\delta}_{t}\triangleq\nabla X^{\delta}_{t}-\nabla X^{x,\xi}_{t},\quad Y^{\delta}_{t}\triangleq\nabla Y^{\delta}_{t}-\nabla Y^{x,\xi}_{t},\quad Z^{\delta}_{t}\triangleq\nabla Z^{\delta}_{t}-\nabla Z^{x,\xi}_{t}.$ (4.10)
Then, we have $(X^{\delta}_{t},Y^{\delta}_{t},Z^{\delta}_{t})$ are uniformly bounded in $L_{r}^{2}$, and they satisfy the following FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{\delta}=\left[H_{xy}^{\delta}X_{t}^{\delta}+H_{yy}^{\delta}Y_{t}^{\delta}+R_{t}^{x}\right]\mathrm{d}t,\\ \begin{aligned} \mathrm{d}Y_{t}^{\delta}&=-\left[H_{xx}^{\delta}X_{t}^{\delta}+H_{xy}^{\delta}Y_{t}^{\delta}-rY_{t}^{\delta}+R_{t}^{y}\right]\mathrm{d}t\\ &\quad+Z_{t}^{\delta}\mathrm{d}B_{t},\end{aligned}\\ X_{0}^{\delta}=0,\end{cases}$ (4.11)
where the remainder terms $R_{t}^{x}$ and $R_{t}^{y}$ are defined as:
$R_{t}^{x}\triangleq$ $\nabla X_{t}^{x,\xi}\left(H_{xy}^{\delta}-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right)$ (4.12)
$+\nabla Y_{t}^{x,\xi}\left(H_{yy}^{\delta}-\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right),$
$R_{t}^{y}\triangleq$ $\nabla X_{t}^{x,\xi}\left(H_{xx}^{\delta}-\partial_{xx}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right)$
$+\nabla Y_{t}^{x,\xi}\left(H_{xy}^{\delta}-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right).$
Applying Itô’s formula to $\mathrm{e}^{-rt}X^{\delta}_{t}Y^{\delta}_{t}$, and noticing that $H_{yy}^{\delta}\leq-\lambda_{1},H_{xx}^{\delta}\geq\lambda_{1},|H_{xy}^{\delta}|\leq\lambda_{2}$, we get that
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left((X_{t}^{\delta})^{2}+(Y_{t}^{\delta})\right)^{2}\mathrm{d}t\right]\leq\frac{2}{r}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\cdot R_{t}^{x}+X_{t}^{\delta}\cdot R_{t}^{y}\right)\mathrm{d}t\right].$ (4.13)
For simplicity, we only estimate the $Y_{t}^{\delta}\cdot\nabla X_{t}^{x,\xi}\left(H_{xx}^{\delta}-\partial_{xx}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right)$ term. Denote
$A_{t}^{\delta}\triangleq\left(H_{xx}^{\delta}-\partial_{xx}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right).$ (4.14)
From the a priori estimate, we know that $Y_{t}^{\delta}$ is uniformly bounded in $L_{r}^{2}$, and $|A_{t}^{\delta}|\leq 2\lambda_{2}$, then there exists a constant $C>0$, such that
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\cdot\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)\mathrm{d}t\right]$ (4.15)
$\leq\left(\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\right)^{2}\mathrm{d}t\right]\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)^{2}\mathrm{d}t\right]\right)^{\frac{1}{2}}$
$\leq C\left(\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)^{2}\mathrm{d}t\right]\right)^{\frac{1}{2}}.$
From the Lipschitz continuity of $\partial_{xx}H$, we know that there exists a constant $\ell>0$, such that
$|A_{t}^{\delta}|\leq\ell\left(|\Delta X^{\delta}_{t}|+|\Delta Y^{\delta}_{t}|\right).$ (4.16)
We consider the following finite measure on $\mathbb{R}_{+}\times\Omega$:
$\mathrm{d}\mathbb{Q}=\mathrm{e}^{-rt}\mathrm{d}t\times\mathrm{d}\mathbb{P}.$ (4.17)
Then, we know
$\int|A^{\delta}|^{2}\mathrm{d}\mathbb{Q}\to 0$ (4.18)
as $\delta\to 0$. This shows that under measure $\mathbb{R}$, $|A^{\delta}|^{2}$ converges in measure to 0. Since $(\nabla X^{x,\xi})^{2}$ is integrable under $\mathbb{Q}$, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ converges in measure to 0. Finally, we know that $(\nabla X^{x,\xi}\cdot A^{\delta})^{2}$ is dominated by an integrable function $4\lambda_{2}^{2}(\nabla X^{x,\xi})^{2}$ under measure Q, and by the Dominated Convergence Theorem, we have
$\lim_{\delta\to 0}\int\left(\nabla X^{x,\xi}\cdot A^{\delta}\right)^{2}\mathrm{d}\mathbb{Q}=0.$ (4.19)
This shows that
$\lim_{\delta\to 0}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(Y_{t}^{\delta}\cdot\nabla X_{t}^{x,\xi}\cdot A_{t}^{\delta}\right)\mathrm{d}t\right]=0.$ (4.20)
Now, we finish the proof.
Now, we consider a function $F(x,\mu,y):\mathbb{R}\times{\cal P}_{2}\times\mathbb{R}\rightarrow\mathbb{R}$ satisfying the following conditions:
  • $F$ is of at most quadratic growth and jointly continuous in $(x,\mu,y)$.
  • $\partial_{x}F(x,\mu,y)$ and $\partial_{y}F(x,\mu,y)$ exist and are Lipschitz continuous in $(x,\mu,y)$.
[…]
We can define a function $U$ on $\mathbb{R}\times{\cal P}_{2}$ by
$U(x,\mu)=\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t\right],$ (4.21)
where $\xi\in\mathbb{L}^{2}({\cal F}_{0};\mu)$. Since we have already established the uniqueness in distribution for the FBSDEs (see [14]), $U$ is well-defined.
We now introduce a lemma about $U$ that will be useful for studying the continuity and differentiability of the value function (1.14).
Lemma 4.3
(i) $\partial_{x}U(x,\mu)$ exists and
[…]
$\qquad+\left.\left.\partial_{y}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\nabla Y^{x,\xi}_{t}\right)\mathrm{d}t\right].$
(ii) $U(x,\mu)$ is jointly continuous in $\mathbb{R}\times{\cal P}_{2}$.
Proof.  (i) We still use the notations in the proof of Theorem 4.2. Fix $x\in\mathbb{R},\xi\in\mathbb{L}^{2}({\cal F}_{0})$ and assume $|\delta|<1$. We have $X_{t}^{x,\xi},{X_{t}^{\xi}},Y_{t}^{x,\xi}$ and $X_{t}^{x+\delta,\xi},Y_{t}^{x+\delta,\xi}$ are all uniformly bounded in $L_{r}^{2}$. From the definition of $U(x,\mu)$, we have
$\frac{U(x+\delta,\mu)-U(x,\mu)}{\delta}=\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(F_{x}^{\delta}\cdot\nabla X^{\delta}_{t}+F_{y}^{\delta}\cdot\nabla Y^{\delta}_{t}\right)\mathrm{d}t\right],$ (4.23)
where
$F_{x}^{\delta}=\int_{0}^{1}\partial_{x}F\left(X_{t}^{x,\xi}+\theta\Delta X^{\delta}_{t},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}+\theta\Delta Y^{\delta}_{t}\right)\mathrm{d}\theta,$ (4.24)
and $F_{y}^{\delta}$ is defined similarly. For simplicity, we let
$\partial_{x}F=\partial_{x}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}),\quad\partial_{y}F=\partial_{y}F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi}).$ (4.25)
Since $\partial_{x}F(x,\mu,y)$ and $\partial_{y}F(x,\mu,y)$ are Lipschitz continuous, we know that
$\left\lVert F_{x}^{\delta}-F_{x}\right\rVert_{r}\to 0,\quad\left\lVert F_{y}^{\delta}-F_{y}\right\rVert_{r}\to 0$ (4.26)
as $\delta\to 0$. Now, we have
$\left\lvert\frac{U(x+\delta,\mu)-U(x,\mu)}{\delta}-\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\partial_{x}F\cdot\nabla X^{x,\xi}_{t}+\partial_{y}F\cdot\nabla Y^{x,\xi}_{t}\right)\mathrm{d}t\right]\right\rvert$ (4.27)
$\leq$ $\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert F_{x}^{\delta}\cdot\nabla X^{\delta}_{t}-\partial_{x}F\cdot\nabla X^{x,\xi}_{t}\right\rvert\mathrm{d}t\right]+\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert F_{y}^{\delta}\cdot\nabla Y^{\delta}_{t}-\partial_{y}F\cdot\nabla Y^{x,\xi}_{t}\right\rvert\mathrm{d}t\right]$
$\triangleq$ $\quad I+II.$
For the first term,
$I\leq$ $\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert F_{x}^{\delta}\cdot\left(\nabla X^{\delta}_{t}-\nabla X^{x,\xi}_{t}\right)\right\rvert\mathrm{d}t\right]+\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left\lvert\left(F_{x}^{\delta}-F_{x}\right)\cdot\nabla X^{x,\xi}_{t}\right\rvert\mathrm{d}t\right]$ (4.28)
$\leq$ $\left\lVert F_{x}^{\delta}\right\rVert_{r}\cdot\left\lVert\nabla X^{\delta}_{t}-\nabla X^{x,\xi}_{t}\right\rVert_{r}+\left\lVert F_{x}^{\delta}-F_{x}\right\rVert_{r}\cdot\left\lVert\nabla X^{x,\xi}_{t}\right\rVert_{r}.$
Since $\left\lVert F_{x}^{\delta}\right\rVert_{r}$ and $\left\lVert\nabla X^{x,\xi}_{t}\right\rVert_{r}$ are uniformly bounded, by Theorem 4.2, we know that I tends to 0 as $\delta\rightarrow 0$. The second term can be treated similarly, then we finish the proof.
(ii) Fix $x\in\mathbb{R},\mu\in{\cal P}_{2}$, and consider another $x^{\prime}\in\mathbb{R},\mu^{\prime}\in{\cal P}_{2}$ such that $|x-x^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})<\delta$. And then take $\xi,\xi^{\prime}\in\mathbb{L}^{2}({\cal F}_{0})$, such that ${\cal L}_{\xi}=\mu,{\cal L}_{\xi^{\prime}}=\mu^{\prime}$. Without loss of generality, we can assum $|\delta|<1$ and $|x-x^{\prime}|+(\mathbb{E}[\xi-\xi^{\prime}]^{2})^{1/2}\leq\delta$ , then all processes appeared in the following proof are uniformly bounded in $L_{r}^{2}$ with respect to $\delta$ . We have
$F(X_{t}^{x^{\prime},\xi^{\prime}},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})$ (4.29)
$=$ $F(X_{t}^{x^{\prime},\xi^{\prime}},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})$
$+F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x^{\prime},\xi^{\prime}})$
$+F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})$
$\triangleq$ $\quad I+II+III.$
For the first term,
$F(X_{t}^{x^{\prime},\xi^{\prime}},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})$ (4.30)
$=$ $\int_{0}^{1}\partial_{x}F(X_{t}^{x,\xi}+\theta(X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi}),{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})\mathrm{d}\theta\cdot(X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi})$
$\triangleq$ $A_{t}\cdot(X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi}).$
Since $\partial_{x}F(x,\mu,y)$ is of at most linear growth, we know that $A_{t}$ is uniformly bounded in $L_{r}^{2}$. By Cauchy’s Inequality and Proposition 3.3, We have
$\mathbb{E}\left[\mathrm{e}^{-rt}|I|\mathrm{d}t\right]\leq\left\lVert A_{t}\right\rVert_{r}\cdot\left\lVert X_{t}^{x^{\prime},\xi^{\prime}}-X_{t}^{x,\xi}\right\rVert_{r}\leq C_{1}\delta.$ (4.31)
For the second term,
$F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi^{\prime}}},Y_{t}^{x^{\prime},\xi^{\prime}})-F(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x^{\prime},\xi^{\prime}})$ (4.32)
$=$ $\int_{0}^{1}\tilde{\mathbb{E}}_{{\cal F}_{t}}\left[\partial_{\mu}F\left(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}+\theta(X_{t}^{\xi^{\prime}}-X_{t}^{\xi})},Y_{t}^{x^{\prime},\xi^{\prime}},{\tilde{X}_{t}^{\xi}+\theta(\tilde{X}_{t}^{\xi^{\prime}}-\tilde{X}_{t}^{\xi})}\right)\cdot\left(\tilde{X}_{t}^{\xi^{\prime}}-\tilde{X}_{t}^{\xi}\right)\right]\mathrm{d}\theta.$
[…]
$\mathbb{E}\left[\mathrm{e}^{-rt}|II|\mathrm{d}t\right]\leq C_{2}\delta.$ (4.33)
The approach for III is identical to that for I. Integrating the analyses of I, II, and III, we obtain there exists a constant $C>0$, such that
$|U(x^{\prime},\mu^{\prime})-U(x,\mu)|\leq C\delta.$ (4.34)
This completes the proof.
4.2 Viscosity solution to elliptic master equation
Theorem 4.4
Under Assumption 3.1 and 4.1, the value function is continuous and satisfies
$\partial_{x}V(x,\mu)=Y_{0}^{x,\xi}.$ (4.35)
Proof.  By the definition of $f(x,\mu,\hat{\alpha}(x,y))$, we have the relationship:
$f(x,\mu,\hat{\alpha}(x,y))=H(x,\mu,y)-\partial_{y}H(x,\mu,y)\cdot y.$ (4.36)
It’s easy to verify that $H(x,\mu,y)$ satisfies all conditions in Lemma 4.3, whereas $\partial_{y}H(x,\mu,y)\cdot y$ fails to do so. We therefore require a more delicate analysis. Fix $x\in\mathbb{R},\xi\in\mathbb{L}^{2}({\cal F}_{0})$, we continue to use the notation from the proof of Theorem 4.2. For any $|\delta|<1$, set
$U^{\delta}\triangleq$ $\frac{1}{\delta}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(\partial_{y}H(X_{t}^{x+\delta,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x+\delta,\xi})\cdot Y^{x+\delta,\xi}_{t}-\partial_{y}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y^{x,\xi}_{t}\right)\mathrm{d}t\right]$ (4.37)
$=$ $\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y^{\delta}_{t}\partial_{y}H(X_{t}^{x+\delta,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x+\delta,\xi})\mathrm{d}t\right]$
$+\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}Y_{t}^{x,\xi}\left(H_{xy}^{\delta}\cdot\nabla X_{t}^{\delta}+H_{yy}^{\delta}\cdot Y_{t}^{\delta}\right)\mathrm{d}t\right]$
and
$U\triangleq$ $\mathbb{E}\bigg[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y_{t}^{x,\xi}\partial_{y}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})$ (4.38)
$+Y_{t}^{x,\xi}\left(\partial_{xy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla X_{t}^{x,\xi}+\partial_{yy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla Y_{t}^{x,\xi}\right)\mathrm{d}t\bigg].$
On the one hand, by an application of Hölder’s inequality, we can obtain that
$\lim_{\delta\to 0}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y^{\delta}_{t}\partial_{y}H(X_{t}^{x+\delta,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x+\delta,\xi})\mathrm{d}t\right]$ (4.39)
$=$ $\mathbb{E}\bigg[\int_{0}^{\infty}\mathrm{e}^{-rt}\nabla Y_{t}^{x,\xi}\partial_{y}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\mathrm{d}t\bigg].$
On the other hand, employing a method similar to that used in the proof of Theorem 4.2, which utilizes the Dominated Convergence Theorem, we can obtain
$\lim_{\delta\to 0}\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}Y_{t}^{x,\xi}\left(H_{xy}^{\delta}\cdot\nabla X_{t}^{\delta}+H_{yy}^{\delta}\cdot Y_{t}^{\delta}\right)\mathrm{d}t\right]$ (4.40)
$=$ $\mathbb{E}\bigg[\int_{0}^{\infty}\mathrm{e}^{-rt}Y_{t}^{x,\xi}\left(\partial_{xy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla X_{t}^{x,\xi}+\partial_{yy}H(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot\nabla Y_{t}^{x,\xi}\right)\mathrm{d}t\bigg].$
Therefore
$\lim_{\delta\to 0}U^{\delta}=U.$ (4.41)
Noting that all the first-order partial derivatives of $f(x,\mu,\hat{\alpha}(x,y))$ have at most linear growth, we obtain from Lemma 4.3 that the value function $V(x,\mu)$ is continuous and $\partial_{x}V(x,\mu)$ exists, which satisfies
$\partial_{x}V(x,\mu)=\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}$ $\left\{\left[\partial_{x}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\right]\cdot\nabla X^{x,\xi}_{t}\right$. (4.42)
$\left.-\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\cdot\nabla Y^{x,\xi}_{t}\right\}\mathrm{d}t.$
Applying Itô’s formula to $\mathrm{e}^{-rt}\nabla X^{x,\xi}_{t}\cdot Y_{t}^{x,\xi}$, we have
$\mathrm{d}\left(\mathrm{e}^{-rt}\nabla X^{x,\xi}_{t}Y_{t}^{x,\xi}\right)$ $=\mathrm{e}^{-rt}\left[-\nabla X^{x,\xi}_{t}\partial_{x}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})+Y_{t}^{x,\xi}\nabla X^{x,\xi}_{t}\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right$. (4.43)
$\quad\left.+Y_{t}^{x,\xi}\nabla Y^{x,\xi}_{t}\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\right]\mathrm{d}t$
$\quad+\mathrm{e}^{-rt}\nabla X^{x,\xi}_{t}Z_{t}^{x}\mathrm{d}B_{t}.$
Note that $\nabla X^{x,\xi}_{0}=1$, and take a sequence $T_{i}\to\infty$ such that
$\mathbb{E}\left[\mathrm{e}^{-rT_{i}}\nabla X^{x,\xi}_{T_{i}}\cdot Y_{T_{i}}^{x,\xi}\right]\to 0.$ (4.44)
Integrating from 0 to $T_{i}$ and taking expectation, after letting $T_{i}\to\infty$ we have:
$Y_{0}^{x,\xi}=\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}$ $\left\{\left[\partial_{x}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})-\partial_{xy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\right]\cdot\nabla X^{x,\xi}_{t}\right$. (4.45)
$\left.-\partial_{yy}H(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}},Y_{t}^{x,\xi})\cdot Y_{t}^{x,\xi}\cdot\nabla Y^{x,\xi}_{t}\right\}\mathrm{d}t.$
Now, we get the desired result.
In preparation for the definition of a viscosity solution of the elliptic master equation, we first define the class of test functions used for that purpose.
Definition 4.5
A function $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ is said to be a test function if the quantities:
[…]
and
[…]
are finite, uniformly in $(x,\mu)$ in any compact subset of $\mathbb{R}\times\mathcal{P}_{2}$.
Now, we present the definition of the viscosity solution for elliptic master equation (1.9).
Definition 4.6
Suppose that $U\in C(\mathbb{R}\times\mathcal{P}_{2})$ and its partial derivative $\partial_{x}U\in C(\mathbb{R}\times\mathcal{P}_{2})$. Then $U$ is called a viscosity subsolution (resp. supersolution) of PDE (1.9) if, for any $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ and any test function $\Psi$, such that $(x^{0},\mu^{0})$ is a local maximum (resp. minimum) of $U-\Psi$, we have
$\begin{split}rU(x^{0},\mu^{0})\leq&H(x^{0},\mu^{0},\partial_{x}U(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\cdot\partial_{y}H(\tilde{\xi}^{0},\mu^{0},\partial_{x}U(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (4.48)
(respectively,
$\begin{split}rU(x^{0},\mu^{0})\geq&H(x^{0},\mu^{0},\partial_{x}U(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\cdot\partial_{y}H(\tilde{\xi}^{0},\mu^{0},\partial_{x}U(\tilde{\xi}^{0},\mu^{0}))\right],\end{split}$ (4.49)
The function $U$ is called a viscosity solution of PDE (1.9) if it is both a viscosity subsolution and a viscosity supersolution.
Theorem 4.7
Under Assumption 3.1 and 4.1, the value function (1.14) is the viscosity solution to the elliptic master equation (1.9).
Proof.  We have proved that $V(x,\mu)$ is a continuous function and $\partial_{x}V(x,\mu)=Y_{0}^{x,\xi}$. Since $\partial_{x}V(x,\mu)$ is also a continuous function, by the flow property, we have
$\partial_{x}V(X_{t}^{x,\xi},{\cal L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi},\quad\partial_{x}V(X_{t}^{\xi},{\cal L}_{X_{t}^{\xi}})=Y_{t}^{\xi}.$ (4.50)
We only show that $V$ is a viscosity subsolution of PDE (1.9). A similar argument will show that it is also a viscosity supersolution of (1.9).
Let $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ be a test function and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $V-\Psi$. It’s natural to get that $\partial_{x}V(x^{0},\mu^{0})=\partial_{x}\Psi(x^{0},\mu^{0})$. We assume without loss of generality that $V(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$. And we suppose that
$\begin{split}r\Psi(x^{0},\mu^{0})>&H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\partial_{y}H(\tilde{\xi},\mu^{0},\partial_{x}V(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (4.51)
Notice that
$\begin{split}&H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))\\ =&f(x^{0},\mu^{0},\hat{\alpha}(x^{0},\partial_{x}V(x^{0},\mu^{0})))+\partial_{x}V(x^{0},\mu^{0})\cdot\partial_{y}H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))\\ =&f(x^{0},\mu^{0},\hat{\alpha}(x^{0},\partial_{x}V(x^{0},\mu^{0})))+\partial_{x}\Psi(x^{0},\mu^{0})\cdot\partial_{y}H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0})).\end{split}$ (4.52)
We can get
$\begin{split}r\Psi(x^{0},\mu^{0})>&f(x^{0},\mu^{0},\hat{\alpha}(x^{0},\partial_{x}V(x^{0},\mu^{0})))+\partial_{x}\Psi(x^{0},\mu^{0})\cdot\partial_{y}H(x^{0},\mu^{0},\partial_{x}V(x^{0},\mu^{0}))\\ &+\frac{1}{2}\partial_{xx}\Psi(x^{0},\mu^{0})\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})+\partial_{\mu}\Psi(x^{0},\mu^{0},\tilde{\xi}^{0})\partial_{y}H(\tilde{\xi},\mu^{0},\partial_{x}V(\tilde{\xi}^{0},\mu^{0}))\right].\end{split}$ (4.53)
It follows from the above that there exists an open subset $O\subset\mathbb{R}\times\mathcal{P}_{2}$ that contains $(x^{0},\mu^{0})$, such that for all $(x,\mu)\in O$,
$\begin{cases}V(x,\mu)\leq\Psi(x,\mu),\\ \begin{aligned} r\Psi(x,\mu)>&f(x,\mu,\hat{\alpha}(x,\partial_{x}V(x,\mu)))+\partial_{x}\Psi(x,\mu)\cdot\partial_{y}H(x,\mu,\partial_{x}V(x,\mu))\\ &+\frac{1}{2}\partial_{xx}\Psi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(x,\mu,\tilde{\xi})+\partial_{\mu}\Psi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}V(\tilde{\xi},\mu))\right].\end{aligned}\end{cases}$ (4.54)
Taking an initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ such that ${\cal L}_{\xi^{0}}=\mu^{0}$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (3.1) and (3.2). We denote $\rho_{t}\triangleq\mathcal{L}_{X_{t}^{\xi^{0}}}$. For some $T>0$, let $\tau$ denote the stopping time
$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\rho_{t})\notin O\}\land T.$ (4.55)
By the flow property, we have that
$\Psi(x^{0},\mu^{0})$ $=V(x^{0},\mu^{0})$ (4.56)
$=\mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t$
$\quad+\mathrm{e}^{-r\tau}\int_{\tau}^{\infty}\mathrm{e}^{-r(t-\tau)}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t\bigg]$
$=\mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t+\mathrm{e}^{-r\tau}V(X_{\tau}^{x^{0},\xi^{0}},\rho_{\tau})\bigg]$
$\leq\mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}f\big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}})\big)\mathrm{d}t+\mathrm{e}^{-r\tau}\Psi(X_{\tau}^{x^{0},\xi^{0}},\rho_{\tau})\bigg].$
By the definition of the test function $\Psi$, we can apply Itô’s formula (2.9) to $\mathrm{e}^{-rt}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$, then we get that
$0\leq$ $\ \mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}\bigg(f\Big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}\big(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}}\big)\Big)$ (4.57)
$\quad-r\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})+\partial_{x}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})\cdot\partial_{y}H(X_{t}^{x^{0},\xi^{0}},\rho_{t},Y_{t}^{x^{0},\xi^{0}})$
$\quad+\frac{1}{2}\partial_{xx}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$
$\quad+\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\Big[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})$
$\qquad+\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})\cdot\partial_{y}H(\tilde{X}_{t}^{\xi^{0}},\rho_{t},\tilde{Y}_{t}^{\xi^{0}})\Big]\bigg)\mathrm{d}t\bigg].$
Then, by the relationship (4.50), we have that
$0\leq$ $\ \mathbb{E}\bigg[\int_{0}^{\tau}\mathrm{e}^{-rt}\bigg(-r\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$ (4.58)
$\quad+f\Big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\hat{\alpha}\big(X_{t}^{x^{0},\xi^{0}},\partial_{x}V(X_{t}^{x^{0},\xi^{0}},\rho_{t})\big)\Big)$
$\quad+\partial_{x}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})\cdot\partial_{y}H\Big(X_{t}^{x^{0},\xi^{0}},\rho_{t},\partial_{x}V(X_{t}^{x^{0},\xi^{0}},\rho_{t})\Big)$
$\quad+\frac{1}{2}\partial_{xx}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t})$
$\quad+\tilde{\mathbb{E}}_{\mathcal{F}_{t}}\bigg[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})$
$\qquad+\partial_{\mu}\Psi(X_{t}^{x^{0},\xi^{0}},\rho_{t},\tilde{X}_{t}^{\xi^{0}})\cdot\partial_{y}H\Big(\tilde{X}_{t}^{\xi^{0}},\rho_{t},\partial_{x}V(\tilde{X}_{t}^{\xi^{0}},\rho_{t})\Big)\bigg]\bigg)\mathrm{d}t\bigg],$
which contradicts (4.54). Now, we finish the proof.
Remark 4.8
For the distribution-independent case, in which $b(x,\mu,a)$ and $f(x,\mu,a)$ reduce to $b(x,a)$ and $f(x,a)$, and consequently the Hamiltonian $H(x,\mu,y)$ becomes $H(x,y)$, we can obtain a stronger conclusion. Considering the unique $L_{r}^{2}$-solution to the reduced mean field FBSDE:
$\begin{cases}\mathrm{d}X_{t}^{x}=\partial_{y}H(X_{t}^{x,\xi},Y_{t}^{x,\xi})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x}=-\left[\partial_{x}{H}(X_{t}^{x,\xi},Y_{t}^{x,\xi})-rY_{t}^{x,\xi}\right]\mathrm{d}t+Z_{t}^{x}\mathrm{d}B_{t},\\ X_{0}^{x}=x,\end{cases}$ (4.59)
we define the value function
$V(x)=\mathbb{E}\bigg[\int_{0}^{+\infty}\mathrm{e}^{-rt}f\big(X_{t}^{x},\hat{\alpha}(X_{t}^{x},Y_{t}^{x})\big)\mathrm{d}t\bigg],$ (4.60)
where $f(x,\hat{\alpha}(x,y))$ can be identified with $H(x,y)-\partial_{y}H(x,y)\cdot y$. From the preceding derivation, it follows that $V(x)$ is a classical solution to the equation
$rU(x)=H(x,\partial_{x}U(x))+\frac{1}{2}\partial_{xx}U(x).$ (4.61)
4.3 Uniqueness of classical solutions to elliptic master equation
Having constructed a viscosity solution to the elliptic master equation via FBSDEs, a natural question arises: is this the unique solution, or at the very least, is the classical solution to the elliptic master equation unique? In [14], we provide an example of an elliptic master equation that admits multiple classical solutions. Considering a case where $r=2,H(x,\mu,y)=2x^{2}-y^{2}$, the PDE (1.9) becomes
$2U(x,\mu)=$ $2x^{2}-\left(\partial_{x}U(x,\mu)\right)^{2}+\frac{1}{2}\partial_{xx}U(x,\mu)$ (4.62)
$+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})-2\partial_{\mu}U(x,\mu,\tilde{\xi})\cdot\partial_{x}U(\tilde{\xi},\mu)\right].$
By solving for the solutions with quadratic form, we obtain the following four solutions:
$U_{1}(x,\mu)=\frac{1}{2}x^{2}+\frac{1}{4},$ (4.63)
$U_{2}(x,\mu)=\frac{1}{2}x^{2}-3x\bar{\mu}+\frac{3}{2}(\bar{\mu})^{2}+\frac{1}{4},$
$U_{3}(x,\mu)=-x^{2}-\frac{1}{2},$
$U_{4}(x,\mu)=-x^{2}+3x\bar{\mu}-\frac{3}{2}(\bar{\mu})^{2}-\frac{1}{2}.$
This demonstrates that the classical solutions to the elliptic master equation are not unique in general. We will now prove that if we consider only the smooth solutions satisfying certain growth and monotonicity conditions, the elliptic master equation admits at most one regular solution.
Definition 4.9
A function $U:\mathbb{R}\times\mathcal{P}_{2}\to\mathbb{R}$ is called a regular solution to the elliptic master equation 1.9 if it satisfies the following conditions:
  1. 1.
    $U\in C^{2,1}$ and $U$ is a classical solution to the elliptic master equation (1.9).
  2. 2.
    $\partial_{x}U\in C^{2,1}$ and $\partial_{x}U$ is a classical solution to PDE (1.8).
  3. 3.
    $U$ is displacement monotone, i.e., for any $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F})$, the following inequality holds:
    $\mathbb{E}\left[(\xi_{1}-\xi_{2})\left(\partial_{x}U(\xi_{1},\mathcal{L}_{\xi_{1}})-\partial_{x}U(\xi_{2},\mathcal{L}_{\xi_{2}})\right)\right]\geq 0.$ (4.64)
  4. 4.
    $U$ is of at most quadratic growth. $\partial_{x}U$ is Lipschitz continuous, and its derivatives, $\partial_{xx}U$, $\partial_{xxx}U$, $\partial_{\mu}\partial_{x}U$ and $\partial_{\tilde{x}\mu}\partial_{x}U$ are of at most linear growth.
Although PDE (1.8) is obtained by differentiating elliptic master equation (1.9) with respect to $x$, we still introduce the second condition. This is because the partial derivatives with respect to $x$ and $\mu$ generally do not commute. By directly requiring $\partial_{x}U$ to be a classical solution to PDE (1.8), we can focus on the main structural issues and avoid secondary technical details. In addition, the requirement that $U$ be displacement monotone stems from the inherent properties of the value function $V(x,\mu)$. From (3.11) and Theorem 4.4, we can obtain
$\mathbb{E}\left[(\xi_{1}-\xi_{2})\left(\partial_{x}V(\xi_{1},\mathcal{L}_{\xi_{1}})-\partial_{x}V(\xi_{2},\mathcal{L}_{\xi_{2}})\right)\right]\geq 0$ (4.65)
for any $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$, which shows that the value function is displacement monotone.
Theorem 4.10
There exists at most one regular solution to the elliptic master equation (1.9).
Proof.  Suppose that $\bar{U}$ is a regular solution. For any $\xi\in\mathbb{L}^{2}({\cal F}_{0})$, consider the following SDE:
$\begin{cases}\mathrm{d}\bar{X}^{\xi}_{t}=\partial_{y}H(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}},\partial_{x}\bar{U}(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}}))\mathrm{d}t+\mathrm{d}B_{t},\\ \bar{X}^{\xi}_{0}=\xi.\end{cases}$ (4.66)
Then, we have, for any $T>0$,
$\mathbb{E}\left[\sup_{t\in[0,T]}|\bar{X}^{\xi}_{t}|^{2}\right]<\infty.$ (4.67)
Let $\bar{Y}^{\xi}_{t}\triangleq\partial_{x}\bar{U}(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}})$ and $\bar{Z}^{\xi}_{t}\triangleq\partial_{xx}\bar{U}(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}})$. Applying Itô’s formula to $\bar{Y}^{\xi}_{t}$ on any finite interval $[0,T]$, we can derive the following BSDE:
$\mathrm{d}\bar{Y}^{\xi}_{t}=-\left[\partial_{x}{H}(\bar{X}^{\xi}_{t},\mathcal{L}_{\bar{X}^{\xi}_{t}},\bar{Y}^{\xi}_{t})-r\bar{Y}^{\xi}_{t}\right]\mathrm{d}t+\bar{Z}^{\xi}_{t}\mathrm{d}B_{t}.$ (4.68)
Thus, we deduce that $(\bar{X}^{\xi},\bar{Y}^{\xi},\bar{Z}^{\xi})$ satisfies the infinite-time FBSDE:
$\begin{cases}\mathrm{d}\bar{X}^{\xi}_{t}=\partial_{y}H(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}},\bar{Y}^{\xi}_{t})\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}\bar{Y}^{\xi}_{t}=-\left[\partial_{x}{H}(\bar{X}^{\xi}_{t},\mathcal{L}_{\bar{X}^{\xi}_{t}},\bar{Y}^{\xi}_{t})-r\bar{Y}^{\xi}_{t}\right]\mathrm{d}t+\bar{Z}^{\xi}_{t}\mathrm{d}B_{t},\\ \bar{X}^{\xi}_{0}=\xi.\end{cases}$ (4.69)
We must emphasize that although the triple $(\bar{X}^{\xi},\bar{Y}^{\xi},\bar{Z}^{\xi})$ satisfies the FBSDE (3.1), it is not necessarily an $L_{r}^{2}$ solution. Consequently, we cannot utilize the uniqueness of solutions for this equation. Let $(X^{\xi},Y^{\xi},Z^{\xi})$ be the unique $L_{r}^{2}$ solution of FBSDE (3.1). We will now utilize the process $(X^{\xi},Y^{\xi},Z^{\xi})$ to prove that the $L_{r}^{2}$-norm of $(\bar{X}^{\xi},\bar{Y}^{\xi},\bar{Z}^{\xi})$ is finite, thereby establishing that $(\bar{X}^{\xi},\bar{Y}^{\xi},\bar{Z}^{\xi})$ is, in fact, equal to $(X^{\xi},Y^{\xi},Z^{\xi})$.
Denote
$I_{t}\triangleq\mathrm{e}^{-rt}(\bar{X}^{\xi}_{t}-X^{\xi}_{t})\left(\partial_{x}\bar{U}(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}})-\partial_{x}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}})\right).$ (4.70)
Since $\bar{U}$ is displacement monotone, we have $\mathbb{E}[I_{T}]\geq 0$, for any $T>0$. To apply Itô’s formula to $I_{t}$, we denote $I_{t}\triangleq N_{t}+M_{t}$, where
$\begin{cases}N_{t}\triangleq\mathrm{e}^{-rt}(\bar{X}^{\xi}_{t}-X^{\xi}_{t})(\bar{Y}^{\xi}_{t}-Y^{\xi}_{t});\\ M_{t}\triangleq\mathrm{e}^{-rt}(\bar{X}^{\xi}_{t}-X^{\xi}_{t})(Y^{\xi}_{t}-\partial_{x}\bar{U}(X^{\xi}_{t},{\cal L}_{X^{\xi}_{t}})).\end{cases}$ (4.71)
Applying Itô’s formula to $N_{t}$ on $[0,T]$ and taking expectation, we derive that
$\mathbb{E}[N_{T}]\leq-\frac{r}{2}\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-rt}\left((\bar{X}^{\xi}_{t}-X^{\xi}_{t})^{2}+(\bar{Y}^{\xi}_{t}-Y^{\xi}_{t})^{2}\right)\mathrm{d}t\right].$ (4.72)
Next, we apply Itô’s formula to $M_{t}$ on $[0,T]$. Noticing that $\partial_{y}H(x,\mu,y)$ is Lipschitz continuous and considering the linear growth condition of the derivatives of $\partial_{x}\bar{U}$, we get that
$\mathbb{E}[M_{T}]\leq$ $\epsilon\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-rt}\left((\bar{X}^{\xi}_{t}-X^{\xi}_{t})^{2}+(\bar{Y}^{\xi}_{t}-Y^{\xi}_{t})^{2}\right)\mathrm{d}t\right]$ (4.73)
$+C(\epsilon)\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-rt}\left(1+(X^{\xi}_{t})^{2}+(Y^{\xi}_{t})^{2}\right)\mathrm{d}t\right].$
Here, $\epsilon$ is a positive constant that is less than $r/2$ and $C(\epsilon)$ depends only on $\epsilon$. Then, we get that there exists a constant $C>0$ (independent of $T$), such that
$\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-rt}\left((\bar{X}^{\xi}_{t}-X^{\xi}_{t})^{2}+(\bar{Y}^{\xi}_{t}-Y^{\xi}_{t})^{2}\right)\mathrm{d}t\right]\leq C\mathbb{E}\left[\int_{0}^{T}\mathrm{e}^{-rt}\left(1+(X^{\xi}_{t})^{2}+(Y^{\xi}_{t})^{2}\right)\mathrm{d}t\right].$ (4.74)
Letting $T\to\infty$, we have that $(\bar{X}^{\xi},\bar{Y}^{\xi},\bar{Z}^{\xi})$ belongs to $L_{r}^{2}$. This means $\partial_{x}\bar{U}(x,\mu)$ is unique.
Now, consider another regular solution $\bar{U}^{\prime}$, we know that $\partial_{x}\bar{U}=\partial_{x}\bar{U}^{\prime}$. Denote $F(\mu)\triangleq\bar{U}(x,\mu)-\bar{U}^{\prime}(x,\mu)$. It satisfies the following PDE:
$\begin{split}rF(\mu)=\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}F(\mu,\tilde{\xi})+\partial_{\mu}F(\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\partial_{x}\bar{U}(\tilde{\xi},\mu))\right].\end{split}$ (4.75)
Applying Itô’s formula to $F({\cal L}_{\bar{X}^{\xi}_{t}})$, we have that
$\mathrm{d}F({\cal L}_{\bar{X}^{\xi}_{t}})$ $={\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}F({\cal L}_{\bar{X}^{\xi}_{t}},\bar{X}^{\xi}_{t})+\partial_{\mu}F({\cal L}_{\bar{X}^{\xi}_{t}},\bar{X}^{\xi}_{t})\partial_{y}H(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}},\partial_{x}\bar{U}(\bar{X}^{\xi}_{t},{\cal L}_{\bar{X}^{\xi}_{t}}))\right]\mathrm{d}t$ (4.76)
$=rF({\cal L}_{\bar{X}^{\xi}_{t}})\mathrm{d}t.$
Thus, we know
$F({\cal L}_{\bar{X}^{\xi}_{t}})=F({\cal L}_{\xi})\mathrm{e}^{rt}.$ (4.77)
Recalling that $F$ is of at most quadratic growth and $\bar{X}^{\xi}\in L_{r}^{2}$, we can choose a sequence $T_{i}\to\infty$ such that $\mathrm{e}^{-rT_{i}}\mathbb{E}[|\bar{X}^{\xi}_{T_{i}}|^{2}]\to 0$. Then we know $F({\cal L}_{\xi})=0$ for any $\xi\in\mathbb{L}^{2}({\cal F}_{0})$. Now, we finish the proof of the uniqueness.
Acknowledgements
Song Y. is financially supported by National Key R&D Program of China (No. 2024YFA1013503 & No. 2020YFA0712700) and the National Natural Science Foundation of China (No. 12431017).